### v1

# LIFT-GS: Cross-Scene Render-Supervised Distillation for 3D Language Grounding

Abstract
Our approach to training 3D vision-language understanding models is to train a feedforward model that makes predictions in 3D, but never requires 3D labels and is supervised only in 2D, using 2D losses and differentiable rendering. The approach is new for vision-language understanding. By treating the reconstruction as a “latent variable”, we can render the outputs without placing unnecessary constraints on the network architecture (e.g. can be used with decoder-only models). For training, only need images and camera pose, and 2D labels. We show that we can even remove the need for 2D labels by using pseudo-labels from pretrained 2D models. We demonstrate this to pretrain a network, and we finetune it for 3D vision-language understanding tasks. We show this approach outperforms baselines/sota for 3D vision-language grounding, and also outperforms other 3D pretraining techniques. Project page: https://liftgs.github.io.
Machine Learning, ICML
1 Introduction
When a user mentions the keys by the door or the blue mug on the table, they use language to indicate a specific set of objects and 3D locations in space. Such 3D language grounding provides a particularly natural interface for people to communicate about their surroundings.
Figure 1: LIFT-GS. We train a 3D vision language grounding model (3D VLG) with point clouds and language inputs by distilling from 2D foundation models without any 3D supervisions.
For AI systems operating in physical spaces, identifying the set of 3D masks or bounding boxes indexed by a language query represents a core functionality; with applications across autonomous navigation, robotic manipulation, and AR/VR. Yet, despite its importance and recent advances in both 2D vision-language grounding and 3D reconstruction, the gap between human and machine performance in 3D Vision-Language Grounding remains one of the most significant challenges in embodied AI.
3D vision-language grounding (3D VLG) is severely constrained by data scarcity. Generative multimodal models are routinely trained on billions to trillions of tokens from language (Achiam et al., 2023; Touvron et al., 2023), images (Radford et al., 2021; Labs, 2023), videos (Polyak et al., 2024; Brooks et al., 2024). Nearer to the task of 3D VLG, models trained for promptable image and video mask segmentation use millions to billions of labeled masks (Kirillov et al., 2023; Ravi et al., 2024). In contrast, existing 3D VLG models are limited to training on only tens of thousands of labeled 3D masks and scenes, restricting their ability.
In this paper, we propose a scalable pretraining approach for 3D VLG that enables grounding language into a set of target 3D masks; without requiring 3D mask labels. Language-Indexed Field Transfer with Gaussian Splatting (LIFT-GS), is supervised on 2D frames directly. It uses differentiable rendering to render the predicted 3D features and masks into 2D frames (e.g. via Gaussian Splatting (Kerbl et al., 2023)). LIFT-GS is then distilled directly from powerful frame-based foundation models, using pseudolabels from SAM (Kirillov et al., 2023), CLIP (Radford et al., 2021), and Llama 3.2 (Meta AI, 2024).
LIFT-GS conducts both Promptable 3D Segmentation and Reconstruction. The render-supervised formulation is highly flexible; as it imposes essentially no constraints on the model input or network architecture. Therefore, in order to demonstrate the effectiveness of differentiable rendering for structured prediction tasks, we choose model inputs and meta-architectures that are scalable and offer advantages for 3D vision-language grounding.
As visual input, LIFT-GS exclusively utilizes point clouds (positions and colors) to align with the preferences for many modern embodied AI applications that operate in real time. Point clouds from SLAM, for example, serve as widely adopted input in robotics and AR/VR. By relying on raw point clouds rather than 2D feature-enhanced point clouds like ConceptFusion (Jatavallabhula et al., 2023b), LIFT-GS eliminates the need for preprocessing and feature fusion during inference, reducing inference time from approximately 60 seconds per frame to just 1 second for the forward pass.
LIFT-GS follows the common MaskFormer (Cheng et al., 2021b) meta-architecture. However, the output tokens from the Encoder and a Mask Decoder are instead interpreted as 3D Gaussians rather than image patches. The entire pipeline is then trained end-to-end by optimizing the 2D loss between the rendered 2D masks and the pseudo-grounding masks.
We validate the effectiveness of LIFT-GS on two popular 3D vision-language grounding (3D VLG) downstream tasks: open-vocabulary 3d instance segmentation and 3D referential grounding. The results show that distillation provides significant improvements over models trained from scratch, and achieves state-of-the-art performance on both tasks. More importantly, the approach consistently improves from both more data and better 2D foundation models, which indicates the potential for further improvement.
Intriguingly, we find that LIFT-GS “pretraining effectively multiplies the fine-tuning dataset” (Hernandez et al., 2021) by a constant factor (roughly 2x), across varying amounts of fine-tuning data. This somewhat counterintuitive observation indeed matches empirical data scaling laws for pretraining in other modalities (Hernandez et al., 2021), where pretraining shifts the power-law scaling in log-dataset size to the left by that constant factor. The fact that the coefficient does not show diminishing returns, even when we use all available fine-tuning data, underscores that 3D VLG models are currently operating in the data scarce regime.
The seeming universality of scaling laws across modalities, and the effectiveness of LIFT-GS pretraining on in 3D VLG settings, suggests that the underlying tools may also be useful at least in other 3D understanding settings. Specifically, render supervised framework can be used with essentially any 3D/4D task or model, provided the results are renderable. This can be combined with knowledge distillation from frame-based models in order to overcome data scarcity in 3D, and LIFT-GS specifically addresses a scarcity of 3D grounded masks. For summary, our contributions are:
  • Differentiable rendering as a tool for training large-scale 3D promptable segmentation models. Instead of proposing a new architecture, the contribution is to train the model using differentiable rendering with a structured 2D grounding loss.
  • A pseudo-labeling strategy for distilling a 2D vision-language grounding pipeline into 3D versions. The pretrained 2D models are needed only for pseudo-labeling and they are not needed during inference.
  • State-of-the-art performance and realistic evaluations. We demonstrate the effectiveness of the approach using sensor pointclouds, common in embodied settings. Rigorous experiments show state-of-the-art performance and reveal scaling properties.
Figure 2: Illustration of 3D Referential Grounding Task. Given an object description, the task requires the model to predict 3D masks for all mentioned noun phrases while ensuring that the located objects satisfy the semantic relationships described in the entire sentence.
2 Related Work
3D Vision-Language Grounding 3D Vision-Language Grounding (3D VLG) refers to the task of mapping language descriptions of objects in a scene, to a set of corresponding 3D masks or bounding boxes that indicate the location of the objects in the observed scene. Despite its importance, there is very little annotated data for 3D VLG available — mainly because there are very little 3D data with annotated masks or bounding boxes period, as well as language annotation for each instance. This provides a major challenge for essentiall all existing models, since they require access to use ground-truth 3D masks or bounding boxes for trainings (Yuan et al., 2021; Roh et al., 2021; Yang et al., 2021; Zhu et al., 2023c, 2024), and many require them during inference in the form of proposal masks or bounding boxes (Fang et al., 2024; Zhang et al., 2023b). This lack of annotated data is a major challenge for all existing models, as the number of 3D scenes with annotated masks or 3D bounding boxes is in the thousands (Dai et al., 2017; Yeshwanth et al., 2023; Somasundaram et al., 2023), compared to the millions of images and videos used for image and video segmentation (Kirillov et al., 2023).
Unlike prior approaches, LIFT-GS introduces a render-supervised approach for 3D VLG. It uses differentiable rendering to train directly on image masks, without requiring 3D bounding boxes or masks during training or inference. This formulation is architecture agnostic and allows the model to be trained using only image and language losses, without the need for 3D masks or bounding boxes. LIFT-GS demonstrates this by training an end-to-end 3D feedforward model. LIFT-GS takes as input a RGB pointcloud and a language utterance as input, and outputs 3D masks.
3D Instance Segmentation 3D instance segmentation refers to the task of predicting masks for a set of instance in a 3D scene, not necessarily given a language prompt. For example, using a predefined set of category labels (Qian et al., 2021; Schult et al., 2022; Hou et al., 2018) or interactive segmentation using input “prompts” (Osep et al., 2024; Zhou et al., 2024; Ma et al., 2024; Chen et al., 2024a).
Because of the lack of 3D data with masks with which to train a 3D segmentation model, a class of recent work has emerged that trains no predictive 3D segmentation model at all, but instead directly lifts powerful pretrained image and video models to 3D using per-scene optimization and multiview geometry. For example, by using depth-unprojection with heuristic view-merging strategies such as voxel-voting (Jatavallabhula et al., 2023a; Zhou et al., 2024; Xu et al., 2023b). Another more recent approach to lift 2D segmentation to 3D is by using differentiable rendering to optimize the 3D representations to minimize the 2D loss between rendered outputs and ground-truth RGB values, 2D features (Kerr et al., 2023; Kim et al., 2024; Dou et al., 2024; Chen et al., 2024b; Gu et al., 2024), or masks (Cen et al., 2023; Xu et al., 2023a). These approaches leverage the large-scale data available for image and video segmentation, but use a fixed per-scene lifting process that does not improve with more training data and takes a long time to run (e.g. minutes per scene on a H100).
These lifted 3D pseudolabels can be used to partially address the lack of 3D data with ground-truth masks, by pre-training a 3D instance segmentation model on the 3D pseudolabels (Genova et al., 2021; Peng et al., 2023). However, this fixed lifting pipeline introduces accumulated errors from inaccuracies in 3D reconstruction and label-merging strategies, creating a bottleneck for overall performance.
LIFT-GS instead trains a promptable 3D segmentation model using only 2D losses, eliminating the need for a fixed lifting process and label-merging strategies. Not relying on heuristic lifting methods or specialized losses (e.g., cross-mask contrastive loss), LIFT-GS is trained on a large-scale dataset using the same mask losses (Dice, Cross-Entropy Focal) as the original frame-level segmentation models. As shown in Appendix Section A.2, LIFT-GS features trained directly on 2D pseudolabels outperform these lifted 3D pseudolabels. And we show effective data scaling in Section 4.4.
Render-Supervised 3D While photometric losses using differentiable rendering (Mildenhall et al., 2020; Kerbl et al., 2023) were originally used for per-scene optimization, render-supervision is emerging as a powerful and generic technique to train predictive 3D models using only 2D losses. For 3D reconstruction, 2D photometric losses can be used to train strong feedforward models (Hong et al., 2024; Tang et al., 2024). PonderV2 (Zhu et al., 2023a) instead uses render-supervision for representation learning, by adding additional constrastive losses on the rendered pixels, using ground-truth category labels to pretrain an Encoder.
LIFT-GS introduces render-supervision for 3D VLG, using mask- and grounding-losses typically used for 2D VLG. LIFT-GS jointly trains an Encoder and Mask Decoder, and for the first time demonstrates using differentiable rendering to train a 3D mask decoder.
3 Method
Figure 3: SAM-CLIP Pseudo-Label Generation. We leverage powerful 2D foundation models to generate pseudo language queries, i.e., CLIP embeddings, along with their corresponding ground-truth 2D masks for training. All pixels within the same mask share the same features.
Figure 4: Architecture Design. LIFT-GS predicts 3D Gaussian Splatting $\mathbf{G}$ and 3D masks $\mathbf{M}$ given a point cloud $\mathbf{P}$ and language query embeddings $\mathbf{Q}$ as inputs. The 3D masks $\mathbf{M}$ are generated by a Transformer-based Mask Decoder.
LIFT-GS is a (pre)training pipeline for 3D vision-language grounding models that uses only 2D supervision. It accomplishes this using differentiable rendering and structured 2D losses, and uses pseudo-labeling for frame annotations.
3.1 Task Formulation
The formulation of 3D Vision-Language Grounding used in LIFT-GS is shown in Figure 2, and based on MDETR (Kamath et al., 2021). The model takes as input a language query, such as “the black chair close to the table near the wall and a pointcloud input. The objective is to predict 3D gaussian masks, for all mentioned noun phrases (chair, table, wall); ensuring that the predicted masks adhere to the semantic relationships described in the language query, via a mask-to-query correspondence matrix.
As shown in Figure 4, LIFT-GS takes as input a point cloud, $\mathbf{P}$, and language query embeddings, $\mathbf{Q}$. It outputs 3D feature Gaussians $\mathbf{G}$ with a correspondence matrix $\mathbf{C}$.
[…]
The pointcloud $\mathbf{P}\in\mathbb{R}^{|P|{\times}6}$ contains positions and RGB colors. LIFT-GS is trained with pointclouds from unprojected posed multi-view RGBD, but any pointcloud reconstruction will work; including pointmap-based methods that do not require depth or camera pose (Wang et al., 2023; Yang et al., 2025; Wang et al., 2025). This makes the method highly flexible, applicable across diverse scenarios, from sparse-view settings to extremely long video sequences.
The Language queries are encoded into embeddings $\mathbf{Q}\in\mathbb{R}^{|Q|{\times}F_{Q}}$, i.e. via CLIP (Radford et al., 2021) in our design. Each embedding can be either from a single word or a full sentence, depending on the desired ganulariry of grounding.
The model predicts 3D Gaussians $\mathbf{G}\in\mathbb{R}^{|G|{\times}|F_{G}|}$ with auxilliary features, enabling the differentiable rendering of 3D information into 2D. In LIFT-GS, the first part of channels of $F_{G}$ are standard Gaussian splatting parameters (i.e. positions $\mu$, covariances $\Sigma$, 3rd-degree SH coefficients $\psi$). The remaining channels are renderable auxiliary 3D attributes: predicted 3D masks $\mathbf{M}\in\mathbb{R}^{|G|{\times}m}$, and renderable auxiliary features: predicted features $\mathbf{F}\in\mathbb{R}^{|G|{\times}F}$.
Following the trend in mask grounding (Liu et al., 2023), the predicted 3D masks $\mathbf{M}\in\mathbb{R}^{|G|\times m}$ contains $m$ predicted masks for each Gaussian, where $m$ is the mask proposal number. Consequently, the rows in the correspondence matrix $\mathbf{C}\in\mathbb{R}^{m\times|Q|}$ represent the log-probability that the $i$-th predicted mask corresponds to each embedding of $\mathbf{Q}$.
3.2 Training 3D with 2D Supervision
Predicting a 3D Gaussian Splatting with desired 3D attributes enables training a 3D model using only 2D supervision, eliminating the need for large-scale 3D annotations. In this section, we first discuss the 2D losses used for training and then describe the process of generating pseudo-labels when perfect labels are unavailable. We show ablations for both losses and pseudo-labels in the experiments.
LIFT-GS utilizes three 2D losses to train the model: $\mathcal{L}_{\text{RGB}}$, $\mathcal{L}_{\text{feat}}$ and $\mathcal{L}_{\text{ground}}$. The rendered image, feature map, and 2D mask are denoted in the equations below as $\tilde{I}\in\mathbb{R}^{H\times W\times 3}$, $\tilde{\mathbf{F}}_{\text{2D}}\in\mathbb{R}^{H\times W\times F}$, and $\tilde{\mathbf{M}}_{\text{2D}}\in\mathbb{R}^{H\times W\times m}$ , respectively. Their corresponding ground-truth counterparts are $I$, $\mathbf{F}_{\text{2D}}$, and $\mathbf{M}_{\text{2D}}$, where $\mathbf{M}_{\text{2D}}\in\mathbb{R}^{H\times W\times K}$.
Reconstruction losses: $\mathcal{L}_{\text{RGB}}$ is the 2D photometric loss, formulated as a combination of $\mathcal{L}_{1}$ loss and SSIM loss (Wang et al., 2004):
$\mathcal{L}_{\text{RGB}}=\lambda_{1}\mathcal{L}_{1}(I,\tilde{I})+\lambda_{2}\mathcal{L}_{\textbf{SSIM}}(I,\tilde{I})$ (2)
Grounding losses: The grounding losses, $\mathcal{L}{\text{mask}}$, measure the prediction quality between the rendered 2D masks, $\tilde{\mathbf{M}}_{\text{2D}}$, and the ground-truth masks, $\mathbf{M}_{\text{2D}}$. Since the predicted and ground-truth masks differ in number, we employ Hungarian matching to compute the set-to-set loss. $\sigma(i)$ is the index of predicted masks matching to $i$-th ground truth mask according to the matching distance $\mathbf{d}_{\text{match}}$.
$\sigma(i)=\arg\min_{j}\mathbf{d}_{\text{match}}(\tilde{\mathbf{M}},\mathbf{M}_{i},\mathbf{C})$ (3)
The mask shape is supervised with $\mathcal{L}_{\text{mask}}$, the combination of Focal (Lin et al., 2017) and Dice (Sudre et al., 2017) loss used in SAM (Kirillov et al., 2023):
$\mathcal{L}_{\text{ground}}\!=\!\frac{1}{K}\sum_{i}^{K}\lambda_{\text{3}}\mathcal{L}_{\text{mask}}(\tilde{\mathbf{M}}_{\text{2D}}^{\sigma(i)},\mathbf{M}_{\text{2D}}^{i})\!+\!\lambda_{\text{4}}\mathcal{L}_{\text{CE}}(\mathbf{C}_{\sigma(i)},i)$ (4)
$\mathcal{L}_{\text{CE}}$ is the correspondence loss, encouraging correct correspondences between predicted masks and language tokens. Specifically, the ground-truth mask set $\mathbf{M}_{\text{2D}}$ consists of $K$ binary masks, each corresponding to a span $\mathbf{Q}_{k}\subseteq\mathbf{Q}$ of language query embeddings $\mathbf{Q}$. $\mathbf{C}_{\sigma(i)}$ represents the correspondence probability of the $\sigma(i)$-th mask proposal across spans of the $|Q|$ language embeddings. We optimize this using a cross-entropy loss.
$\mathcal{L}_{\text{CE}}(\mathcal{C}_{\sigma(i)},i)=-\log\frac{\exp(\mathbf{C}_{\sigma(i),i})}{\sum_{j}^{K}\exp(\mathbf{C}_{\sigma(i),j})}$ (5)
Feature loss: $\mathcal{L}_{\text{feat}}$ is the feature rendering loss which measures the difference between rendered feature map $\tilde{\mathbf{F}}_{\text{2D}}$ and ground-truth feature map $\mathbf{F}_{\text{2D}}$. As in (Zhu et al., 2023b), LIFT-GS uses a contrastive loss.
$\mathcal{L}_{\text{feat}}=\frac{1}{H\times W}\sum_{u,v}^{H,W}-\log\frac{\exp(\tilde{\mathbf{f}}_{(u,v)}\cdot\mathbf{f}_{k})}{\sum_{j}\exp(\tilde{\mathbf{f}}_{(u,v)}\cdot\mathbf{f}_{j})}$ (6)
$\tilde{\mathbf{f}}{(u,v)}$ is the feature vector of $\tilde{\mathbf{F}}_{\text{2D}}$ at location $(u,v)$. ${\mathbf{f}}_{j}$ denotes a batch of unique feature vectors from $\mathbf{F}$, and $\mathbf{f}k$ is the ground-truth feature vector corresponding to $\tilde{\mathbf{f}}{(u,v)}$.
3.3 SAM-CLIP 2D Pseudo-Label
Although LIFT-GS doesn’t need 3D annotations, getting high-quality 2D supervision remains challenging. We show how leveraging 2D foundation models for pseudo-label generation enables reasonable zero-shot performance and significantly enhances downstream tasks after fine-tuning.
As shown in Figure 3, we leverage 2D foundation models to generate pseudo-labels, including pseudo language queries and corresponding 2D masks. For each image, we apply SAM (Kirillov et al., 2023) to obtain segmentation masks. For each segmented region, we extract a CLIP (Radford et al., 2021) image embedding as the pseudo language query embedding. Since CLIP’s text and image embeddings share the same feature space, LIFT-GS can take text embeddings as input during inference. During pretraining, we concatenate these CLIP embeddings to form $\mathbf{Q}$ and construct $\mathbf{C}$, using the corresponding 2D masks as ground truth.
Figure 5: Zero-Shot 3D Segmentation. Trained using only 2D pseudo-labels, LIFT-GS can localize objects in 3D from real text inputs in a zero-shot manner. From left to right, we visualize the input point clouds, segmented 3D masks(in yellow), rendered images from predicted 3DGS, and rendered segmentation masks. Language queries include both high-level abstract concepts (e.g., white) and detailed descriptions (e.g., black cabinet near the wall).
Trained with 2D pseudo-labels, LIFT-GS can perform zero-shot 3D grounding using real text queries without fine-tuning, as shown in Figure 5. However, the zero-shot model suffers from low accuracy and struggles with complex expressions, a common limitation of CLIP-based methods that function as bag-of-words models (Yuksekgonul et al., 2023). Future improvements in pseudo-labeling, such as captioning (Meta AI, 2024) and 2D language grounding models (Liu et al., 2023), could alleviate the need for fine-tuning altogether. In this work, we focus on fine-tuning.
Overall, LIFT-GS shows that even simple pseudo-labeling strategies can be effectively distilled into 3D models. Our experiments show that pretraining with 2D pseudo-labels substantially boost performance on downstream tasks.
3.4 Architecture
LIFT-GS  is network-agnostic, imposing few architectural constraints and being readily adapted to other architectures. We describe the network used in experiments, as in Figure 4.
Encoder: Predicting 3DGS from Point Cloud
Encoder takes point clouds as input and predicts 3D Gaussian splatting (without masks) and per-point features as latent variables. Generally, the number of predicted Gaussians $|G|$ can differ from the number of input points $|P|$.
However, our focus is on demonstrating the effectiveness of differentiable rendering for structured prediction tasks like Promptable 3D segmentation. So in our experiments, we set $|G|=|P|$ with a bijective mapping to ensure consistency with point cloud-based evaluation tasks. For comparison to prior work (Zhu et al., 2023b) , LIFT-GS uses a SparseConv UNet (Contributors, 2022) as the Encoder Backbone to get per-point features, and uses an MLP as Gaussian Head to regress Gaussian parameters of $\mathbf{G}$ with features $\mathbf{F}$.
Decoder: Mask Decoder for 3D VLG
The Decoder takes per-point features from the Encoder and language embeddings $\mathbf{Q}$ as input to predict the 3D mask $\mathbf{M}\in\mathbb{R}^{N\times m}$ and the correspondence matrix $\mathbf{C}\in\mathbb{R}^{m\times|Q|}$. The predicted $\mathbf{M}$ serves as a renderable feature of the 3DGS $\mathbf{G}$ and is subsequently used to generate 2D masks. While this work focuses on 3D vision-language grounding (3D VLG), the proposed framework is a general pipeline that can be extended to other 3D tasks with renderable outputs.
We adopt a 3D variant of the Transformer-based mask decoder in MaskFormer (Cheng et al., 2021a; Jain et al., 2025), inspired by the success of MDETR (Kamath et al., 2021) and BUTD (Jain et al., 2021). As illustrated in Figure 4, the 3D Mask Decoder Transformer takes three inputs: visual tokens from per-point features, language tokens from $\mathbf{Q}$, and $m$ learnable mask proposal tokens. It outputs $m$ predicted mask embeddings and a correspondence matrix $\mathbf{C}$. The 3D masks $\mathbf{M}$ are computed as the dot product between visual tokens and output mask embeddings. These similarity scores are then differentiably rendered into mask images $\tilde{\mathbf{M}}_{\text{2D}}$.
4 Experiments
Table 1: Open-Vocabulary 3D Instance Segmentation. We evaluate our model on ScanNet200 by using category names as text queries and compare it against SOTA models.
[…]
OpenScene (Peng et al., 2023) 11.7 17.8 15.2
OpenMask3D (Takmaz et al., 2023) 15.4 23.1 19.9
PQ3D (Zhu et al., 2024) 20.2 32.5 28.0
LIFT-GS-Scratch 22.5 35.1 30.7
LIFT-GS 25.7 40.2 35.0
$\Delta$ $+3.2\uparrow$ $+5.1\uparrow$ $+4.3\uparrow$
In this section, we first provide training details and demonstrate how pretraining significantly enhances downstream task performance through a set of carefully designed ablations. Additionally, we uncover several intriguing findings by scaling pretraining and fine-tuning data, as well as exploring the impact of different 2D foundation models.
4.1 Training Details
We provide details below with more details in Appendix.
Training Datasets We use ScanNet (Dai et al., 2017) as the primary dataset for downstream task fine-tuning and evaluation, as its annotations form the basis for established benchmarks. Our method’s flexibility enables training on diverse 3D datasets without requiring language annotations or instance masks. For pretraining, we primarily use ScanNet (Dai et al., 2017) and ScanNet++(Yeshwanth et al., 2023) and also explore Taskonomy(Zamir et al., 2018) and Aria Synthetic (Somasundaram et al., 2023).
Architecture Our pipeline imposes no constraints on the 3D architecture or inductive biases, allowing flexibility in model selection. For experiments, we use SparseConv UNet as the encoder backbone, a widely adopted architecture for point cloud processing. The Mask Decoder is an 8-layer Transformer decoder with a hidden state size of 512.
Training Details We train the Encoder and Decoder jointly using a batch size of 32 for 76k steps on 32 A100 GPUs. We use AdamW (Loshchilov & Hutter, 2017) optimizer with a learning rate of 1e-4 and weight decay of 1e-4.
4.2 Evaluation on 3D Vision-Language Grounding
We fine-tune and evaluate our pretrained model on two representative 3D VLG tasks: 3D open-vocabulary instance segmentation and 3D referential grounding. Our results, shown in Tables 1 and 2, demonstrate significant improvements over models trained from scratch and achieve state-of-the-art performance with pertaining.
4.2.1 Grounding Simple Nouns in 3D
We first evaluate simple grounding for simple noun-phrases, using object categories without spatial relationships. Following the protocol in (Zhu et al., 2024), we convert the standard 3D instance segmentation benchmark on ScanNet into an open-vocabulary 3D instance segmentation task. The categories of objects are used as language queries, which are input to the model to predict the corresponding 3D masks.
Evaluation setting: We evaluate using the standard metric mAP, a measure of mask overlap averaged across categories. We fine-tune LIFT-GS for 500 epochs.
Results: Compared against the state-of-the-art baselines PQ3D (Zhu et al., 2024) and OpenMask3D (Takmaz et al., 2023), our pretrained model (LIFT-GS) achieves substantial performance gains (mAP 25.7% vs 20.2%), as shown in Table 1. It significantly outperforms its counterpart trained from scratch (LIFT-GS-Scratch mAP +3.2%).
Table 2: 3D Referential Grounding. We report top-1 accuracy with various IoU thresholds (0.25, 0.5).
SR3D NR3D ScanRefer
Method Acc@25 Acc@50 Acc@25 Acc@50 Acc@25 Acc@50
Mesh PC
LanguageRefer (Roh et al., 2021) 39.5 - 28.6 - - -
SAT-2D (Yang et al., 2021) 35.4 - 31.7 - 44.5 30.1
BUTD-DETR (Jain et al., 2021) 52.1 - 43.3 - 52.2 39.8
3D-VisTA (Zhu et al., 2023c) 56.5 51.5 47.7 42.2 51.0 46.2
PQ3D (Zhu et al., 2024) 62.0 55.9 52.2 45.0 56.7 51.8
Sensor PC + Bounding Box Proposals using Mesh PC
3D-VisTA (Zhu et al., 2023c) 47.2 43.2 42.1 37.4 46.4 42.5
Sensor PC
BUTD-DETR (Jain et al., 2021) 43.3 28.9 32.2 19.4 42.2 27.9
LIFT-GS-Scratch 44.0 28.8 37.2 23.1 45.0 29.5
LIFT-GS 50.9 36.5 43.7 29.7 49.7 36.4
$\Delta$ +6.9(16%) +7.7(27%) +6.5(17%) +6.6(29%) +4.7(10%) +6.9(23%)
Table 3: Comparison with other Pretraining Baseline. LIFT-GS clearly outperforms Ponder-v2 and its variant Ponder-v2$\dagger$, which is trained on the same SAM-CLIP features as ours.
Model Acc@@@@0.25 Acc@@@@0.5 Acc@@@@0.75
Scratch 42.19 27.23 9.66
Ponder-v2 (official) 40.92 25.97 8.84
Ponder-v2$\dagger$ 45.40 29.36 9.29
LIFT-GS 47.53 33.75 13.49
4.2.2 Grounding Complex Phrases in 3D
Next, we examine grounding multiple objects using more complex phrases that contain spatial references, referred to as 3D Referential Grounding (3D RG).
Evaluation Setting. We evaluate LIFT-GS on the most common 3D Referential Grounding benchmarks: ScanRefer (Chen et al., 2019), SR3D, and NR3D (Achlioptas et al., 2020; Abdelreheem et al., 2022). We use standard top-1 accuracy as the evaluation metric, considering a predicted bounding box correct if its IoU with the ground truth exceeds 0.25 or 0.5. Since LIFT-GS outputs masks instead of axis-aligned bounding boxes, we derive bounding boxes by extracting the extreme corner points from the point cloud within the predicted masks.
LIFT-GS is designed to be practical. To ensure a more realistic evaluation for embodied applications, we introduce two modifications: (1) we predict 3D masks without assuming the availability of ground-truth 3D bounding boxes, and (2) we utilize sensor point clouds (Sensor PC) from RGB-D scans instead of mesh-derived point clouds (Mesh PC).
Existing methods are typically trained on mesh-derived point clouds (Mesh PC), requiring creating scene meshes from RGB-D scans, processing, and sampling point clouds. They generally rely on segments derived from clustering centers generated during human annotation. Since such information is unavailable in real-world applications, we instead use raw sensor point clouds (Sensor PC) from RGB-D scans. This setting is inherently more challenging, as reflected in the significant performance drop of BUTD-DETR (Jain et al., 2021) when transitioning from Mesh PC to Sensor PC (Table 2), consistent with findings in (Jain et al., 2024).
Baselines We compare LIFT-GS against the state-of-the-art two-stage methods, 3D-VisTA (Zhu et al., 2023c) and PQ3D (Zhu et al., 2024), as well as the SOTA single-stage method, BUTD-DETR (Jain et al., 2021). All two-stage baselines assume access to ground-truth 3D masks or boxes during inference, so we re-evaluate them using predicted boxes from the SOTA object detector Mask3D (Schult et al., 2022). For fairness, we re-train 3D-VisTA and BUTD-DETR on sensor point clouds. Because PQ3D uses multiple backbones and a multi-stage training pipeline, we were not able to reproduce PQ3D on the sensor point cloud setting.
Results Our model without pretraining (LIFT-GS-Scratch) achieves slightly better performance than the state-of-the-art single-stage method BUTD-DETR (Jain et al., 2021), likely due to architectural similarities with extra modifications.
With pretraining, LIFT-GS achieves significant improvements across all three datasets, with relative gains of $10\%-30\%$, demonstrating the effectiveness of our pretraining approach. Notably, LIFT-GS outperforms 3D-VisTA in Acc@$25$, despite 3D-VisTA being a two-stage method with bounding box proposals from Mask3D using Mesh PC.
4.3 Ablation and Analyses
We conduct an in-depth analysis of the proposed method through a series of ablation and scaling experiments. For these evaluations, we use a model pretrained only on ScanNet as the baseline. To simplify the presentation for the ablations, we report results on the combined evaluation set of ScanRefer, SR3D, and NR3D. Additionally, we report the higher accuracy threshold Acc@$0.75$.
Compare to SOTA pretraining methods We compare to Ponder-v2 (Zhu et al., 2023b), a state-of-the-art method for point cloud pretraining. Ponder-v2 takes point clouds as input and predicts voxel grids, which are then used to render 2D images via NeRF (specifically NeuS, which additionally models surfaces (Wang et al., 2021)). Pretraining is supervised by 2D photometric loss and CLIP features, where features are computed based on per-pixel text labels for the mask (GT category labels and masks for each instance).
We found that the official Ponder-v2 checkpoint fails to improve performance on 3D referential grounding, likely due to the limited size and diversity of its per-pixel text labels. Therefore, we retrain the Ponder-v2 while using the SAM-CLIP pseudolabels described in Section 3.3. Training on the pseudolabels (indicated by Ponder-v2$\dagger$ in Table 3) significantly improves Acc@$0.25$ +4.5% vs. the official Ponder-v2, and +3.2% vs. training from scratch. Since both approaches use render-supervision, this experiment underscores the advantages of pseudolabel distillation in data-scarce regimes. Moreover our method does not rely on human-annotated per-pixel text labels, making it highly scalable to other datasets, as shown in Table 6.
Table 4: Loss Ablation. We show the impact of different pretraining losses on 3D referential grounding task. $\mathcal{L}_{\text{ground}}$ significantly improves results, particularly at high IoU thresholds.
Model $\mathcal{L}_{\text{ground}}$ $\mathcal{L}_{\text{RGB}}$ $\mathcal{L}_{\text{feat}}$ Acc@@@@0.25 Acc@@@@0.5 Acc@@@@0.75
Scratch 42.19 27.23 9.66
- 46.34 31.54 12.50
- 46.67 31.81 12.45
- 47.69 31.35 11.36
- 47.53 33.75 13.49
Loss Ablation Existing pretraining pipelines primarily focus on the encoder (Zhu et al., 2023b; Banani et al., 2021), whereas the render-supervised formulation can pretrain the entire architecture in a unified manner. To assess the impact of the grounding loss, which applies only to the decoder, we compare models pretrained with different losses in Table 4.
A model trained with $\mathcal{L}_{\text{ground}}$ alone (row 2) still substantially improves downstream task performance, performing only slightly worse than the model trained with the full loss (row 5), demonstrating the effectiveness of the grounding loss. Furthermore, comparing models with and without $\mathcal{L}{\text{ground}}$ (row 5 vs. row 4) clearly shows that $\mathcal{L}_{\text{ground}}$ significantly enhances downstream performance, particularly in more challenging scenarios (IoU thresholds of 0.5 and 0.75).
Table 5: Fine-tune Data Scaling. We show Acc@$0.5$ results with different ratio of fine-tuning data on referential grounding task.
Finetuning Data Ratio 10% 20% 50% 100%
Scratch 6.93 15.04 23.00 27.23
LIFT-GS 14.70 23.03 28.89 33.75
4.4 Data Scaling
Finetuning Data Scaling We limit the fine-tuning data size to $0.1$, $0.2$, and $0.5$ of the full dataset for 3D referential grounding to analyze performance variation with different fine-tuning ratios. The results are presented in Figure 6 and Table 5, with additional findings provided in Appendix A.3.
Intriguingly, we find that LIFT-GS “pretraining effectively multiplies the fine-tuning dataset” (Hernandez et al., 2021) by a constant factor (roughly 2x), across varying amounts of fine-tuning data. For instance, the pretrained model fine-tuned on $20\%$ of the data achieves performance comparable to training from scratch with $50\%$ of the data. The benefits of pretraining are even more pronounced at higher IoU thresholds. Overall, a pretrained model can achieve the same performance as a model trained from scratch on all the data using only $30{-}40\%$ of the fine-tuning data.
This observation matches empirical data scaling laws for pretraining in other modalities (Hernandez et al., 2021). And interestingly, the fact that the coefficient does not show diminishing returns, even when we use all available fine-tuning data, which underscores that 3D VLG models are currently operating in the data scarce regime.
Figure 6: Fine-tune Data Scaling. We show how Grounding Accuracy changes with increasing Data Ratio from 0.1 to 1.0.
Pretraining Data Scaling A key advantage of our framework is its ability to pretrain on large-scale RGB-D scans without requiring 3D or language annotations, ensuring scalability. We scale the pretraining dataset and analyze its impact on fine-tuning performance, as shown in Table 6.
[…]
Since our method doesn’t need 3D masks or language annotations, expanding the pretraining dataset with real-world scans is fairly easy. We anticipate continuous improvements as larger-scale data collections become available.
Table 6: Pretraining on OOD data. Adding more pretraining data from ScanNet++ improves performance. Taskonomy and Arial helped less than ScanNet++, likely due to distribution difference.
Pretraining Data Acc@@@@0.25 Acc@@@@0.5 Acc@@@@0.75
Scannet 47.53 33.75 13.49
+Scannet++ 48.29 34.35 14.06
++ Taskonomy and Arial 48.49 34.41 14.35
4.5 2D Foundation Models Scaling and Exploration
Our pipeline leverages powerful 2D foundation models to generate pseudo-labels. Here, we investigate their impact by analyzing performance variations with different 2D foundation models, with results presented in Table 7.
Weaker CLIP and SAM The main experiments use SAM-H and CLIP-L for pseudo-labeling. Replacing them with smaller models, MobileSAM (Zhang et al., 2023a)(ViT-tiny)  and CLIP-B, leads to a noticeable performance drop, especially at higher accuracy thresholds. This suggests that our model benefits from advancements in 2D foundation models, highlighting the importance of stronger 2D models.
Table 7: 2D Foundation Model Exploration.
2D Models Acc@0.25 Acc@0.5 Acc@0.75
SAM-B + CLIP-B 46.31 31.50 12.41
SAM-H + CLIP-L 47.53 33.75 13.49
SAM-H + LLAMA-Caption 47.50 32.78 13.25
Captions from VLMs We explore an alternative pseudo-labeling strategy by combining SAM with the LLAMA-V model (details in Appendix). After segmenting objects in 2D images using SAM, we prompt LLAMA-V to describe the segmented regions and encode these descriptions into CLIP embeddings.
Pretraining with this pseudo-labeling method achieves performance comparable to our original pipeline with SAM-H and CLIP-L, highlighting the flexibility and extensibility of our framework. We believe text-based captions hold significant potential for future research, and our pipeline is well-positioned to benefit from advancements in this area.
5 Conclusion
We present LIFT-GS, a model-agnostic pipeline for training 3D VLG models without 3D supervision by leveraging Gaussian Splatting and 2D foundation models. Beyond enabling zero-shot 3D grounding, LIFT-GS significantly improves downstream performance when fine-tuned with limited 3D data. It also scales effectively, benefiting from larger datasets and stronger 2D foundation models, suggesting continuous improvements with increased data. Limitations and future directions are discussed in the Appendix.
Appendix A More Details
A.1 Training Details
LIFT-GStakes point clouds and posed RGB images for training. For efficiency, we preprocess point clouds and posed RGB images, caching the processed features.
Point clouds originate from multi-frame RGB-D scans. We unproject them using depth information and fuse the unprojections into the final point clouds. Each dataset sample is preprocessed into 5cm-resolution point cloud chunks with corresponding posed RGB images.
For 2D pseudo-labels, we precompute SAM-CLIP features and cache them. Given the large size of the feature map, we decompose it into two components: Semantics and Index2Semantics.
Semantics: A tensor of shape $H\times W$, where each pixel stores the index of the segment it belongs to. Index2Semantics: A tensor of shape $N\times F$, where $N$ is the number of unique segments, and $F$ is the CLIP feature dimension. This decomposition significantly reduces storage costs. When computing the feature rendering loss $\mathcal{L}_{\text{feat}}$, we directly use features from Index2Semantics for contrastive loss.
Each training sample consists of a sparse point cloud and a posed image with corresponding SAM-CLIP features. We randomly sample up to 8 unique instances, using their CLIP features as pseudo language queries and their masks as target 2D masks. To ensure mask quality, we filter out masks smaller than 1024 pixels.
Randomly sampling instances is important for training, especially for zero-shot segmentation, as it prevents the model to reconstruct the whole images given all the input embeddings.
For grounding loss, we assign weights of 15.0, 2.0, and 6.0 to the mask cross-entropy loss, soft token loss, and Dice loss, respectively. We also use a photometric loss (L1 and SSIM) with a weight of 1.0 and a feature loss with a weight of 0.1.
UNet Encoder: 8 layers, maximum channel dimension of 256, output feature dimension of 96. MaskDecoder: 8-layer Transformer decoder with a hidden state size of 512. It uses 256 learnable mask proposal tokens, generating 256 masks. Each Transformer block has 8 attention heads, a feedforward MLP of dimension 2048, and a dropout ratio of 0.15. Language Encoder: We use clip-vit-large-patch14, with a feature dimension of 768.
A.2 Comparison to 3D pseudolabels
Table 8: Comparison to 3D pseudolabels. A mask decoder trained on top of frozen LIFT-GS features matches and even outperforms a decoder trained on top of lifted 3D pseudolabels (voxel-pooled ConceptFusion (Jatavallabhula et al., 2023a)). LIFT-GS learns to pool features in 3D in order to optimally reproduce the pseudolabels after rendering, which outperforms using a hand-crafted aggregation. Note: in this experiment we used a more expressive mask decoder in this experiment with a larger MLP ratio, which improves the results for all methods, including LIFT-GS.
Features Acc@@@@0.25 Acc@@@@0.5
Scratch (RGB) 44.1 30.6
3D pseudolabels 50.1 34.7
2D pseudolabels (LIFT-GS features) 51.8 38.3
LIFT-GS (finetuned) 54.7 40.5
A.3 Data Scaling Results
A similar trend is observed for 3D open-vocabulary instance segmentation, though the benefits of pretraining are slightly less pronounced due to the task’s lower complexity. This aligns with our findings that pretraining is more beneficial for challenging tasks, such as those with higher IoU thresholds or greater complexity.
A.4 VLM Captions
Figure 7: Finetunning Data Scaling on Open Vocabulary 3D Instance Segmentation. We show how mAP changes along with increasing Data Ratio from 0.1 to 1.0
We explore using vision-language models (VLMs) to generate captions for each SAM-segmented object and encode these captions into CLIP embeddings as pseudo language queries.
Specifically, given a SAM-segmented region, we draw a red bounding box on the 2D image and highlight the masked region using alpha blending, as shown in Figure 8. We then prompt a VLM, such as LLama-3.2v, with the following instruction:
You are a helpful assistant for image captioning. You are given an image with a red bounding box specifying the object of interest. Caption that object in a few words, keeping it precise and concise. The object is also slightly highlighted. Examples output: ”a red traffic light,” ”the box near the wall.” Just output the caption; no other text is needed.
This approach leverages VLM-generated textual descriptions to improve pseudo-language queries for training.
Appendix B Discussion and Limitations
Figure 8: Input Image to VLM for Captions. Given the segments from SAM, we draw red bounding box around the segments and ask VLM models to describe the segments inside the red bounding boxes.
The core contribution of LIFT-GS is training a 3D model without 3D supervision by leveraging differentiable rendering and distilling knowledge from 2D foundation models. This approach is novel and motivated by the fact that 2D foundation models, trained on vast amounts of 2D data, currently outperform any existing 3D model. Distilling knowledge from these powerful 2D models presents a promising and scalable direction for 3D learning.
Our proposed pipeline is general and unified. Beyond 3D masks, any renderable 3D attributes can, in principle, be trained using 2D supervision. This idea could extend to dynamic scenes and other properties, opening new opportunities for 3D model training.
However, LIFT-GS is inherently constrained by how well we leverage 2D foundation models for pseudo-labeling. Currently, we use CLIP image embeddings as text queries, but CLIP’s claim of a shared embedding space for images and text is imperfect. In practice, these embeddings can differ significantly, leading to challenges in zero-shot 3D segmentation.
Our CLIP-SAM features may not be optimal pseudo-labels for pretraining, and we anticipate that improved pseudo-labeling strategies will lead to better scaling properties, stronger performance, and even robust zero-shot 3D segmentation without fine-tuning. Addressing our current limitations presents a key opportunity for future work.
Although LIFT-GS significantly improves performance and surpasses the single-stage SOTA method BUTD-DETR (Jain et al., 2021), it still falls short of two-stage SOTA methods like 3D-VisTA (Zhu et al., 2024) on 3D referential grounding at an IoU threshold of 0.5. A robust single-stage 3D VLG model would have a major impact across various applications. We hope that our architecture-agnostic pretraining pipeline can further enhance future models.

### v2

@@
-LIFT-GS: Cross-Scene Render-Supervised Distillation for 3D Language Grounding
+From Thousands to Billions: 3D Visual Language Grounding via Render-Supervised Distillation from 2D VLMs
 Abstract
-Our approach to training 3D vision-language understanding models is to train a feedforward model that makes predictions in 3D, but never requires 3D labels and is supervised only in 2D, using 2D losses and differentiable rendering.
-The approach is new for vision-language understanding.
-By treating the reconstruction as a “latent variable”, we can render the outputs without placing unnecessary constraints on the network architecture (e.g.
-can be used with decoder-only models).
-For training, only need images and camera pose, and 2D labels.
-We show that we can even remove the need for 2D labels by using pseudo-labels from pretrained 2D models.
-We demonstrate this to pretrain a network, and we finetune it for 3D vision-language understanding tasks.
-We show this approach outperforms baselines/sota for 3D vision-language grounding, and also outperforms other 3D pretraining techniques.
+3D vision-language grounding faces a fundamental data bottleneck: while 2D models train on billions of images, 3D models have access to only thousands of labeled scenes–a six-order-of-magnitude gap that severely limits performance.
+We introduce LIFT-GS, a practical distillation technique that overcomes this limitation by using differentiable rendering to bridge 3D and 2D supervision.
+LIFT-GS predicts 3D Gaussian representations from point clouds and uses them to render predicted language-conditioned 3D masks into 2D views, enabling supervision from 2D foundation models (SAM, CLIP, LLaMA) without requiring any 3D annotations.
+This render-supervised formulation enables end-to-end training of complete encoder-decoder architectures and is inherently model-agnostic.
+LIFT-GS achieves state-of-the-art results with 25.7% mAP on open-vocabulary instance segmentation (vs.
+20.2% prior SOTA) and consistent 10-30% improvements on referential grounding tasks.
+Remarkably, pretraining effectively multiplies fine-tuning datasets by 2×, demonstrating strong scaling properties that suggest 3D VLG currently operates in a severely data-scarce regime.
 Project page: https://liftgs.github.io.
@@
 Such 3D language grounding provides a particularly natural interface for people to communicate about their surroundings.
-Figure 1: LIFT-GS. We train a 3D vision language grounding model (3D VLG) with point clouds and language inputs by distilling from 2D foundation models without any 3D supervisions.
-For AI systems operating in physical spaces, identifying the set of 3D masks or bounding boxes indexed by a language query represents a core functionality; with applications across autonomous navigation, robotic manipulation, and AR/VR.
-Yet, despite its importance and recent advances in both 2D vision-language grounding and 3D reconstruction, the gap between human and machine performance in 3D Vision-Language Grounding remains one of the most significant challenges in embodied AI.
-3D vision-language grounding (3D VLG) is severely constrained by data scarcity.
-Generative multimodal models are routinely trained on billions to trillions of tokens from language (Achiam et al., 2023; Touvron et al., 2023), images (Radford et al., 2021; Labs, 2023), videos (Polyak et al., 2024; Brooks et al., 2024).
-Nearer to the task of 3D VLG, models trained for promptable image and video mask segmentation use millions to billions of labeled masks (Kirillov et al., 2023; Ravi et al., 2024).
-In contrast, existing 3D VLG models are limited to training on only tens of thousands of labeled 3D masks and scenes, restricting their ability.
-In this paper, we propose a scalable pretraining approach for 3D VLG that enables grounding language into a set of target 3D masks; without requiring 3D mask labels.
-Language-Indexed Field Transfer with Gaussian Splatting (LIFT-GS), is supervised on 2D frames directly.
-It uses differentiable rendering to render the predicted 3D features and masks into 2D frames (e.g.
-via Gaussian Splatting (Kerbl et al., 2023)).
-LIFT-GS is then distilled directly from powerful frame-based foundation models, using pseudolabels from SAM (Kirillov et al., 2023), CLIP (Radford et al., 2021), and Llama 3.2 (Meta AI, 2024).
-LIFT-GS conducts both Promptable 3D Segmentation and Reconstruction.
-The render-supervised formulation is highly flexible; as it imposes essentially no constraints on the model input or network architecture.
-Therefore, in order to demonstrate the effectiveness of differentiable rendering for structured prediction tasks, we choose model inputs and meta-architectures that are scalable and offer advantages for 3D vision-language grounding.
-As visual input, LIFT-GS exclusively utilizes point clouds (positions and colors) to align with the preferences for many modern embodied AI applications that operate in real time.
-Point clouds from SLAM, for example, serve as widely adopted input in robotics and AR/VR.
-By relying on raw point clouds rather than 2D feature-enhanced point clouds like ConceptFusion (Jatavallabhula et al., 2023b), LIFT-GS eliminates the need for preprocessing and feature fusion during inference, reducing inference time from approximately 60 seconds per frame to just 1 second for the forward pass.
-LIFT-GS follows the common MaskFormer (Cheng et al., 2021b) meta-architecture.
-However, the output tokens from the Encoder and a Mask Decoder are instead interpreted as 3D Gaussians rather than image patches.
-The entire pipeline is then trained end-to-end by optimizing the 2D loss between the rendered 2D masks and the pseudo-grounding masks.
-We validate the effectiveness of LIFT-GS on two popular 3D vision-language grounding (3D VLG) downstream tasks: open-vocabulary 3d instance segmentation and 3D referential grounding.
-The results show that distillation provides significant improvements over models trained from scratch, and achieves state-of-the-art performance on both tasks.
-More importantly, the approach consistently improves from both more data and better 2D foundation models, which indicates the potential for further improvement.
-Intriguingly, we find that LIFT-GS “pretraining effectively multiplies the fine-tuning dataset” (Hernandez et al., 2021) by a constant factor (roughly 2x), across varying amounts of fine-tuning data.
-This somewhat counterintuitive observation indeed matches empirical data scaling laws for pretraining in other modalities (Hernandez et al., 2021), where pretraining shifts the power-law scaling in log-dataset size to the left by that constant factor.
-The fact that the coefficient does not show diminishing returns, even when we use all available fine-tuning data, underscores that 3D VLG models are currently operating in the data scarce regime.
-The seeming universality of scaling laws across modalities, and the effectiveness of LIFT-GS pretraining on in 3D VLG settings, suggests that the underlying tools may also be useful at least in other 3D understanding settings.
-Specifically, render supervised framework can be used with essentially any 3D/4D task or model, provided the results are renderable.
-This can be combined with knowledge distillation from frame-based models in order to overcome data scarcity in 3D, and LIFT-GS specifically addresses a scarcity of 3D grounded masks.
-For summary, our contributions are:
-  • Differentiable rendering as a tool for training large-scale 3D promptable segmentation models.
-Instead of proposing a new architecture, the contribution is to train the model using differentiable rendering with a structured 2D grounding loss.
-  • A pseudo-labeling strategy for distilling a 2D vision-language grounding pipeline into 3D versions.
-The pretrained 2D models are needed only for pseudo-labeling and they are not needed during inference.
-  • State-of-the-art performance and realistic evaluations.
-We demonstrate the effectiveness of the approach using sensor pointclouds, common in embodied settings.
-Rigorous experiments show state-of-the-art performance and reveal scaling properties.
-Figure 2: Illustration of 3D Referential Grounding Task.
-Given an object description, the task requires the model to predict 3D masks for all mentioned noun phrases while ensuring that the located objects satisfy the semantic relationships described in the entire sentence.
+For AI systems operating in physical spaces, identifying the 3D masks or bounding boxes indexed by language queries represents a core functionality, with applications across autonomous navigation, robotic manipulation, and AR/VR.
+Figure 1: LIFT-GS  Overview.
+We train a powerful 3D vision language grounding model (i.e., 3D mask decoder) with point clouds and language as inputs by learning from 2D VLM foundation models without any 3D supervision.
+Yet despite its importance, 3D vision-language grounding (3D VLG) faces a fundamental bottleneck: data scarcity.
+While 2D vision-language models are trained on billions of labeled images and masks (Achiam et al., 2023; Touvron et al., 2023; Radford et al., 2021; Labs, 2023), existing 3D VLG models have access to only thousands of labeled 3D scenes and masks.
+This six-order-of-magnitude gap in data availability severely limits the capabilities of current 3D grounding systems, creating one of the most significant challenges in embodied AI.
+A common workaround to this scarcity constructs 3D feature fields from 2D features (e.g., CLIP embeddings) and performs text queries via dot products between the text and 3D embeddings.
+Although this provides good generalization, performance degrades with more detailed descriptions typical of real-world queries, as illustrated in Figure 3.
+From this perspective, the dual-encoder approach falls short of 3D grounding as it contradicts a core grounding requirement.
+In this paper, we ask: can we combine the best part of both pipelines, i.e., training a powerful grounding model while still overcoming the data scarcity by learning from powerful 2D models?
+The key insight is that differentiable rendering provides a natural bridge between 3D and 2D.
+If we can predict 3D masks and render them into 2D views, we can supervise them using 2D foundation models that have been trained on internet-scale data.
+This approach could enable training 3D models without any 3D mask annotations.
+We introduce Language-Indexed Field Transfer with Gaussian Splatting (LIFT-GS), which implements this idea as a practical training pipeline.
+Given a point cloud and language query, LIFT-GS predicts 3D Gaussian representations that can be rendered into multiple 2D views.
+These rendered masks are then supervised using pseudo-labels from 2D foundation models.
+We show results where SAM provides mask supervision (Kirillov et al., 2023), and CLIP or LLaMa provide language understanding (Radford et al., 2021; Meta AI, 2024).
+The approach effectively distills internet-scale 2D knowledge into 3D understanding.
+This render-supervised formulation offers several key advantages.
+First, it is inherently architecture-agnostic; specifying only the outputs leaves flexibility in underlying model design.
+Second, this allows us to overcome fundamental scaling limitations by training a large transformer decoder instead of previous dual-encoder approaches (as shown in Fig 3(Zhu et al., 2023b; Gu et al., 2024).
+Third, the approach is highly practical: LIFT-GS operates directly on raw point clouds from sensors, such as the outputs from SLAM or SfM systems, eliminating the preprocessing and feature fusion required by methods like ConceptFusion (Jatavallabhula et al., 2023b; Arnaud et al., 2025), reducing inference time from  60 seconds to just 1 second.
+Our experiments validate the effectiveness of this approach.
+LIFT-GS achieves state-of-the-art performance on standard 3D VLG benchmarks, with 25.7% mAP on open-vocabulary instance segmentation (vs.
+20.2% previous SOTA) and consistent 10-30% relative improvements on referential grounding tasks.
+More importantly, we observe that across data scales for SFT, pretraining effectively ”multiplies” the fine-tuning dataset by approximately a constant factor (2x).
+That is, a pretrained model with 50% of fine-tuning data matches the performance of training from scratch with 100% data.
+This somewhat counterintuitive observation indeed matches empirical data scaling laws for pretraining in other modalities (Hernandez et al., 2021), and the fact that this scaling coefficient remains constant without diminishing returns across data scales adds an empirical data point that 3D VLG currently operates in the very low-data regime 111(Hernandez et al., 2021) define a “low-data regime as having 10% or less of the amount of data it would take to get to 99% of the performance that infinite data would yield.”.
+The implications extend beyond 3D grounding.
+Render supervision is powerful, but we demonstrate that it can serve as a bridge for large-scale knowledge transfer from 2D foundation models to 3D models.
+Any 3D/4D task with renderable outputs can potentially leverage 2D supervision.
+As 2D foundation models continue to improve and scale, 3D models trained using render-supervised distillation are positioned to benefit.
+This opens the possibility of training 3D understanding models at the scale of 2D datasets–which would represent a fundamental shift from the current paradigm of limited 3D annotations.
+To summarize, our contributions are:
+  • A render-supervised training pipeline for 3D vision-language grounding that requires only 2D supervision.
+We show how differentiable rendering enables training 3D models with 2D losses, eliminating dependence on scarce 3D annotations.
+  • Demonstrating a pseudo-labeling strategy for distilling 2D foundation models into 3D. LIFT-GS shows using SAM, CLIP, and LLMs to generate 2D supervision.
+  • State-of-the-art performance in realistic evaluations. LIFT-GS achieves SOTA results using sensor point clouds common in embodied settings, with detailed ablations revealing scaling properties.
+Figure 2: 3D Referential Grounding. For each mentioned instance in a text description, predict a 3D mask and map it to corresponding text tokens.
 2 Related Work
-3D Vision-Language Grounding 3D Vision-Language Grounding (3D VLG) refers to the task of mapping language descriptions of objects in a scene, to a set of corresponding 3D masks or bounding boxes that indicate the location of the objects in the observed scene.
-Despite its importance, there is very little annotated data for 3D VLG available — mainly because there are very little 3D data with annotated masks or bounding boxes period, as well as language annotation for each instance.
-This provides a major challenge for essentiall all existing models, since they require access to use ground-truth 3D masks or bounding boxes for trainings (Yuan et al., 2021; Roh et al., 2021; Yang et al., 2021; Zhu et al., 2023c, 2024), and many require them during inference in the form of proposal masks or bounding boxes (Fang et al., 2024; Zhang et al., 2023b).
-This lack of annotated data is a major challenge for all existing models, as the number of 3D scenes with annotated masks or 3D bounding boxes is in the thousands (Dai et al., 2017; Yeshwanth et al., 2023; Somasundaram et al., 2023), compared to the millions of images and videos used for image and video segmentation (Kirillov et al., 2023).
-Unlike prior approaches, LIFT-GS introduces a render-supervised approach for 3D VLG.
-It uses differentiable rendering to train directly on image masks, without requiring 3D bounding boxes or masks during training or inference.
-This formulation is architecture agnostic and allows the model to be trained using only image and language losses, without the need for 3D masks or bounding boxes.
-LIFT-GS demonstrates this by training an end-to-end 3D feedforward model.
-LIFT-GS takes as input a RGB pointcloud and a language utterance as input, and outputs 3D masks.
-3D Instance Segmentation 3D instance segmentation refers to the task of predicting masks for a set of instance in a 3D scene, not necessarily given a language prompt.
-For example, using a predefined set of category labels (Qian et al., 2021; Schult et al., 2022; Hou et al., 2018) or interactive segmentation using input “prompts” (Osep et al., 2024; Zhou et al., 2024; Ma et al., 2024; Chen et al., 2024a).
-Because of the lack of 3D data with masks with which to train a 3D segmentation model, a class of recent work has emerged that trains no predictive 3D segmentation model at all, but instead directly lifts powerful pretrained image and video models to 3D using per-scene optimization and multiview geometry.
-For example, by using depth-unprojection with heuristic view-merging strategies such as voxel-voting (Jatavallabhula et al., 2023a; Zhou et al., 2024; Xu et al., 2023b).
-Another more recent approach to lift 2D segmentation to 3D is by using differentiable rendering to optimize the 3D representations to minimize the 2D loss between rendered outputs and ground-truth RGB values, 2D features (Kerr et al., 2023; Kim et al., 2024; Dou et al., 2024; Chen et al., 2024b; Gu et al., 2024), or masks (Cen et al., 2023; Xu et al., 2023a).
-These approaches leverage the large-scale data available for image and video segmentation, but use a fixed per-scene lifting process that does not improve with more training data and takes a long time to run (e.g.
-minutes per scene on a H100).
-These lifted 3D pseudolabels can be used to partially address the lack of 3D data with ground-truth masks, by pre-training a 3D instance segmentation model on the 3D pseudolabels (Genova et al., 2021; Peng et al., 2023).
-However, this fixed lifting pipeline introduces accumulated errors from inaccuracies in 3D reconstruction and label-merging strategies, creating a bottleneck for overall performance.
-LIFT-GS instead trains a promptable 3D segmentation model using only 2D losses, eliminating the need for a fixed lifting process and label-merging strategies.
-Not relying on heuristic lifting methods or specialized losses (e.g., cross-mask contrastive loss), LIFT-GS is trained on a large-scale dataset using the same mask losses (Dice, Cross-Entropy Focal) as the original frame-level segmentation models.
-As shown in Appendix Section A.2, LIFT-GS features trained directly on 2D pseudolabels outperform these lifted 3D pseudolabels.
-And we show effective data scaling in Section 4.4.
-Render-Supervised 3D While photometric losses using differentiable rendering (Mildenhall et al., 2020; Kerbl et al., 2023) were originally used for per-scene optimization, render-supervision is emerging as a powerful and generic technique to train predictive 3D models using only 2D losses.
-For 3D reconstruction, 2D photometric losses can be used to train strong feedforward models (Hong et al., 2024; Tang et al., 2024).
-PonderV2 (Zhu et al., 2023a) instead uses render-supervision for representation learning, by adding additional constrastive losses on the rendered pixels, using ground-truth category labels to pretrain an Encoder.
-LIFT-GS introduces render-supervision for 3D VLG, using mask- and grounding-losses typically used for 2D VLG.
-LIFT-GS jointly trains an Encoder and Mask Decoder, and for the first time demonstrates using differentiable rendering to train a 3D mask decoder.
+2.1 The Data Scarcity Challenge in 3D VLG
+3D Vision-Language Grounding (3D VLG) maps language descriptions to corresponding 3D masks or bounding boxes in observed scenes (Yuan et al., 2021; Roh et al., 2021; Yang et al., 2021).
+Despite its fundamental importance for embodied AI, existing 3D VLG datasets contain only thousands of annotated scenes (Dai et al., 2017; Yeshwanth et al., 2023) compared to billions of images used for training large multimodal models (Meta AI, 2024).
+This scarcity stems from prohibitive annotation costs.
+Creating 3D instance masks requires minutes per example even with assisted tools (Dai et al., 2017), and human verification remains necessary (Arnaud et al., 2025; Majumdar et al., 2024).
+All existing 3D VLG methods require ground-truth 3D masks or bounding boxes during training (Yuan et al., 2021; Zhu et al., 2023c, 2024; Zhang et al., 2024), with many also requiring them during inference (Fang et al., 2024; Zhang et al., 2023b).
+This creates a fundamental scaling barrier–these approaches cannot leverage the vast 2D data that has driven progress in 2D understanding.
+2.2 Bridging 2D and 3D: From Lifting to Learning
+2D-to-3D Lifting via Optimization.
+Recent work addresses data scarcity by lifting 2D models to 3D through per-scene optimization.
+Methods use depth unprojection with heuristic merging (e.g., voxel voting) (Jatavallabhula et al., 2023a; Zhou et al., 2024; Xu et al., 2023b) or differentiable rendering to optimize 3D representations matching 2D features (Kerr et al., 2023; Kim et al., 2024; Gu et al., 2024) or masks (Cen et al., 2023; Xu et al., 2023a).
+While these leverage 2D foundation models, they suffer from: (1) slow optimization (minutes per scene), (2) accumulated errors from reconstruction and merging, and (3) inability to improve with more data.
+The fixed lifting pipelines may be a bottleneck that explains why LIFT-GS outperforms models trained on these lifted pseudolabels (Genova et al., 2021; Peng et al., 2023).
+Render-Supervised Learning.
+Differentiable rendering enables training 3D models directly by rendering their predictions into 2D and supervising via 2D losses.
+While initially used for reconstruction (Hong et al., 2024; Tang et al., 2024; Cao et al., 2024; Szymanowicz et al., 2025),  (Irshad et al., 2024; Zhu et al., 2023b) use it to representation learning, like PonderV2 adds CLIP losses on rendered pixels (Zhu et al., 2023a).
+However, existing 3D VLG methods have significant limitations: PonderV2 only trains encoders and relies on ground-truth category labels, while other methods apply simple photometric losses without leveraging 2D VLMs.
+LIFT-GS overcomes these challenges by extending render-supervision in two key ways: (1) it enables training unified encoder-decoder architectures for 3D VLG tasks, and (2) it performs knowledge distillation from 2D foundation models (VLMs, SAM, CLIP) through pseudo-labeling, removing the need for 3D annotations during pretraining.
+2.3 Architecture: From Dot-Products to Joint Attention
+Figure 3: 3D grounding with CLIP-style (dual-decoder) method.
+Grounding heatmaps from a representative approach (Guo et al., 2024).
+Heatmaps are computed using dot product similarity between visual tokens and text tokens (as in the CLIP objective), encoded independently.
+This performs effectively with very short prompts, but fails with more detailed queries, as shown in the image.
+LIFT-GS addresses this by jointly predicting tokens using a transformer decoder with expressive attention masks (see Figure 6 and experiments).
+Limitations of Dual-Encoder Methods.
+Almost all prior 3D grounding approaches that distill from 2D vision-language models eventually compute the mask as the dot-product similarity between 3D features and text embeddings (Radford et al., 2021; Guo et al., 2024; Qin et al., 2023; Kerr et al., 2023; Peng et al., 2023).
+When modality embeddings are computed independently, as is the case in CLIP, these “dual-encoder” models behave as bag-of-words systems (Yuksekgonul et al., 2022).
+Using these encoders causes 3D models to inherit these fundamental limitations; as shown in the limited ability of 3D dual-encoder models to handle relational language crucial for referential grounding (e.g., ”the chair next to the table”) (Fig.
+3).
+Multimodal Decoders.
+Transformer decoders address the bag-of-words behavior by jointly processing the modalities together through learned attention mechanisms that enable proper handling of spatial relationships and multi-object references.
+Following large multimodal language models (LLaMA 3, GPT-4o, and Qwen 2.5), as well as recent 3D VLG SotA (Kamath et al., 2021; Jain et al., 2025; Arnaud et al., 2025), LIFT-GS employs a decoder-based architecture.
+LIFT-GS introduces the grounding loss described in Sec.
+3.2 to train the decoder.
+2.4 Foundation Model Distillation at Scale
+Recent work on scaling up pseudolabeling pipelines shows that although 2D foundation models are increasingly capable (Hong et al., 2023; Arnaud et al., 2025), they currently exhibit significant limitations in spatial understanding over multiple frames, frequently hallucinating 3D spatial relationships (Majumdar et al., 2024; Yang et al., 2024).
+This is why LIFT-GS uses pseudolabels for pretraining: just as LLMs require supervised fine-tuning (SFT) to align noisy internet text with desired behaviors, LIFT-GS leverages noisy pseudolabels for large-scale pretraining, then uses 3D VLG SFT for state-of-the-art performance.
+Our scaling analysis reveals suggests that even imperfect 2D spatial understanding generates meaningful training signal (multiplying fine-tuning data effectiveness by 2× in our experiments).
+As 2D models advance in spatial reasoning, this transfer benefit should amplify, potentially reducing reliance on 3D annotations and moving toward more zero-shot spatial understanding.
+This positions our approach at the intersection of three key trends: (1) the shift from optimization-based to learning-based 3D understanding, (2) the move from dual-encoder to multimodal decoder architectures for complex language grounding, and (3) the emergence of foundation model distillation as a solution to 3D data scarcity.
+Our experiments demonstrate that this approach not only achieves state-of-the-art performance but also exhibits strong scaling properties, suggesting significant potential for using cross-scene render-supervised distillation with improved multimodal foundation models and in other settings in besides 3D VLG.
 3 Method
@@
 The 3D masks $\mathbf{M}$ are generated by a Transformer-based Mask Decoder.
-LIFT-GS is a (pre)training pipeline for 3D vision-language grounding models that uses only 2D supervision.
-It accomplishes this using differentiable rendering and structured 2D losses, and uses pseudo-labeling for frame annotations.
+LIFT-GS is a (pre-)training pipeline for 3D VLG without 3D GT annotation.
+During training, it renders the predicted 3D masks from the target viewpoints for 2D supervision; during testing, 3D masks are used as outputs.
+This section covers the 3D VLG formalism in Sec.3.1, 2D loss used for supervision in Sec.3.2, pseudolabel generation in Sec.3.4, and implementation details in Sec.3.3.
 3.1 Task Formulation
-The formulation of 3D Vision-Language Grounding used in LIFT-GS is shown in Figure 2, and based on MDETR (Kamath et al., 2021).
-The model takes as input a language query, such as “the black chair close to the table near the wall and a pointcloud input.
-The objective is to predict 3D gaussian masks, for all mentioned noun phrases (chair, table, wall); ensuring that the predicted masks adhere to the semantic relationships described in the language query, via a mask-to-query correspondence matrix.
-As shown in Figure 4, LIFT-GS takes as input a point cloud, $\mathbf{P}$, and language query embeddings, $\mathbf{Q}$.
-It outputs 3D feature Gaussians $\mathbf{G}$ with a correspondence matrix $\mathbf{C}$.
+3.1.1 3D Vision-Language Grounding
+Figure 2 shows an example using a point cloud input ($\mathbf{P})$ and the text query ($\mathbf{Q}$) “the black chair close to the table near the wall”.
+Following MDETR (Kamath et al., 2021) and (Jain et al., 2025), LIFT-GS outputs a set of ($m=256$) 3D mask candidates $\mathbf{M}$, and the correspondence matrix $\mathbf{C}$
 […]
-The pointcloud $\mathbf{P}\in\mathbb{R}^{|P|{\times}6}$ contains positions and RGB colors.
-LIFT-GS is trained with pointclouds from unprojected posed multi-view RGBD, but any pointcloud reconstruction will work; including pointmap-based methods that do not require depth or camera pose (Wang et al., 2023; Yang et al., 2025; Wang et al., 2025).
-This makes the method highly flexible, applicable across diverse scenarios, from sparse-view settings to extremely long video sequences.
-The Language queries are encoded into embeddings $\mathbf{Q}\in\mathbb{R}^{|Q|{\times}F_{Q}}$, i.e.
-via CLIP (Radford et al., 2021) in our design.
-Each embedding can be either from a single word or a full sentence, depending on the desired ganulariry of grounding.
-The model predicts 3D Gaussians $\mathbf{G}\in\mathbb{R}^{|G|{\times}|F_{G}|}$ with auxilliary features, enabling the differentiable rendering of 3D information into 2D.
-In LIFT-GS, the first part of channels of $F_{G}$ are standard Gaussian splatting parameters (i.e.
-positions $\mu$, covariances $\Sigma$, 3rd-degree SH coefficients $\psi$).
-The remaining channels are renderable auxiliary 3D attributes: predicted 3D masks $\mathbf{M}\in\mathbb{R}^{|G|{\times}m}$, and renderable auxiliary features: predicted features $\mathbf{F}\in\mathbb{R}^{|G|{\times}F}$.
-Following the trend in mask grounding (Liu et al., 2023), the predicted 3D masks $\mathbf{M}\in\mathbb{R}^{|G|\times m}$ contains $m$ predicted masks for each Gaussian, where $m$ is the mask proposal number.
-Consequently, the rows in the correspondence matrix $\mathbf{C}\in\mathbb{R}^{m\times|Q|}$ represent the log-probability that the $i$-th predicted mask corresponds to each embedding of $\mathbf{Q}$.
-3.2 Training 3D with 2D Supervision
-Predicting a 3D Gaussian Splatting with desired 3D attributes enables training a 3D model using only 2D supervision, eliminating the need for large-scale 3D annotations.
-In this section, we first discuss the 2D losses used for training and then describe the process of generating pseudo-labels when perfect labels are unavailable.
-We show ablations for both losses and pseudo-labels in the experiments.
-LIFT-GS utilizes three 2D losses to train the model: $\mathcal{L}_{\text{RGB}}$, $\mathcal{L}_{\text{feat}}$ and $\mathcal{L}_{\text{ground}}$.
-The rendered image, feature map, and 2D mask are denoted in the equations below as $\tilde{I}\in\mathbb{R}^{H\times W\times 3}$, $\tilde{\mathbf{F}}_{\text{2D}}\in\mathbb{R}^{H\times W\times F}$, and $\tilde{\mathbf{M}}_{\text{2D}}\in\mathbb{R}^{H\times W\times m}$ , respectively.
-Their corresponding ground-truth counterparts are $I$, $\mathbf{F}_{\text{2D}}$, and $\mathbf{M}_{\text{2D}}$, where $\mathbf{M}_{\text{2D}}\in\mathbb{R}^{H\times W\times K}$.
-Reconstruction losses: $\mathcal{L}_{\text{RGB}}$ is the 2D photometric loss, formulated as a combination of $\mathcal{L}_{1}$ loss and SSIM loss (Wang et al., 2004):
+Matrix $\mathbf{C}\in\mathbb{R}^{m{\times}|Q|}$: indicates the correspondence (i.e.
+probability logits) between $m$ 3D mask candidates and $|Q|$ text tokens, enabling the mapping of each text query to its most probable mask candidate based on the logits.
+3D Mask $\mathbf{M}\in\mathbb{R}^{m{\times}N}$: stores mask logits over $N$ Gaussian primitives for $m$ 3D mask candidates.
+Each logit $\mathbf{M}_{i,j}$ represents the probability (after applying sigmoid) that the $j$-th Gaussian primitive is included in the $i$-th mask candidate.
+This dual design elegantly handles two key challenges: (1) text tokens can refer to multiple instances (e.g., “the chairs”), and (2) instances can be referenced multiple times with different descriptions throughout the text (e.g.
+both ”chair” and ”it”).
+It also extends readily to segmenting other modalities by adding more mappings.
+The pointcloud has $|P|$ points and is of shape $|P|\times 6$ (XYZ + RGB), and the query token embeddings and are a matrix of size: $\mathbf{Q}\in\mathbb{R}^{|Q|{\times}F_{Q}}$.
+The text mapping $\mathbf{C}$ and Gaussian/point cloud mapping $\mathbf{M}$ are shown, with each instances highlighted in a different color in Figure 2.
+3.1.2 3D VLG with Gaussian Masks $\mathbf{M}$
+Existing 3D VLG methods are limited by costly point cloud annotation, where annotation for point cloud masks is by far the costliest and slowest step in 3D VLG data collection.
+This is because human annotators must carefully segment the mask using a brush or assisted tool, which takes on the order of minutes per example (Dai et al., 2017).
+Even with model assistance, human verification is required in practice (Arnaud et al., 2025; Majumdar et al., 2024).
+This restricts training to only thousands of scenes, making data scarcity the key bottleneck.
+LIFT-GS addresses this by leveraging differentiable rendering, distilling knowledge from multimodal vision-language models trained on billions of images.
+It predicts the 3D Gaussians from the point cloud input feed-forwardly, which are used for rendering later.
+3D Gaussians are represented with xyz locations $\mathbb{R}^{N\times 3}$, covariance matrices $\mathbb{R}^{N\times 6}$, color matrices $\mathbb{R}^{N\times 3}$ and feature embeddings $\mathbb{R}^{N\times F}$.
+LIFT-GS predicts $\mathbf{G}\in\mathbb{R}^{N\times(m+12+512)}$ containing $m$ masks plus 12 channels for shape/location/color and 512 for feature loss:
+[…]
+In practice, training the masks using differentiable rendering adds only a small overhead during training.
+With the shapes of the inputs and outputs specified, the sections below describe each component: the losses used, pseudolabel generation, and the model architecture.
+3.2 Losses
+With differentiable rendering, LIFG-GS enables training 3D VLG models using the simplest 2D grounding losses.
+During training, LIFT-GS only requires (sparse) point cloud, 2D posed images as inputs without any other 3D annotations, which data can be easily obtained from RGB-D videos or SfM (Yang et al., 2025; Wang et al., 2025a, 2023, b; Leroy et al., 2024) as done in  (Szymanowicz et al., 2025).
+LIFT-GS utilizes two groups of losses to train the model: $\mathcal{L}_{\text{ground}}$ for grounding (Jain et al., 2025), and per-pixel losses $\mathcal{L}_{\text{PP}}$ commonly used to improve results with differentiable rendering.
+For reference, rendered 2D masks, rgb images, and feature maps are denoted as $\tilde{\mathbf{M}}_{\text{2D}}\in\mathbb{R}^{H\times W\times m}$, $\tilde{\mathbf{F}}_{\text{2D}}\in\mathbb{R}^{H\times W\times F}$, and $\tilde{I}\in\mathbb{R}^{H\times W\times 3}$, respectively.
+Their corresponding ground-truth counterparts have analogous shapes: $\mathbf{M}_{\text{2D}}$, $\mathbf{F}_{\text{2D}}$, and $I$, where $\mathbf{M}_{\text{2D}}\in\mathbb{R}^{H\times W\times K}$.
+3.2.1 Grounding losses:
+LIFT-GS uses the MDETR-style mask grounding loss: $\mathcal{L}_{\text{CE}}$ and $\mathcal{L}_{\text{mask}}$ with $\mathbf{d}_{\text{match}}$ (matching distance).
+Since the number of predicted and ground-truth masks may differ, we apply Hungarian matching to pair them based on $\mathbf{d}_{\text{match}}$, and then optimize the matched predictions using $\mathcal{L}_{\text{ground}}$.
+We use the 2D variant of these losses under 2D rendering supervision, which can be easily replaced by their 3D counterpart (on 3D Gaussian centers) when 3D labels are available.
+$\mathcal{L}_{\text{ground}}\!=\!\frac{1}{K}\sum_{i}^{K}\lambda_{\text{3}}\mathcal{L}_{\text{mask}}(\tilde{\mathbf{M}}_{\text{2D}}^{\sigma(i)},\mathbf{M}_{\text{2D}}^{i})\!+\!\lambda_{\text{4}}\mathcal{L}_{\text{CE}}(\mathbf{C}_{\sigma(i)},i)$ (3)
+$\sigma(i)=\arg\min_{j}\mathbf{d}_{\text{match}}(\tilde{\mathbf{M}},\mathbf{M}_{i},\mathbf{C})$ (4)
+Cross Entropy Loss, $\mathcal{L}_{\text{CE}}$: this loss supervises the correspondence between 3D mask candidates and input language tokens (i.e., the matrix $\mathbf{C}$) by framing it as a (soft) classification problem.
+Recent work suggests that using BCE may produce sharper masks for long VLG text queries (Jain et al., 2025; Arnaud et al., 2025; Zhai et al., 2023), echoing similar findings in image segmentation (Cheng et al., 2021a).
+$\mathcal{L}_{\text{CE}}(\mathcal{C}_{\sigma(i)},i)=-\log\frac{\exp(\mathbf{C}_{\sigma(i),i})}{\sum_{j}^{K}\exp(\mathbf{C}_{\sigma(i),j})}$ (5)
+Mask loss, $\mathcal{L}_{\text{mask}}$: this loss supervises the predicted 3D masks by comparing them to paired ground-truth 3D masks.
+Following SAM (Kirillov et al., 2023), we apply a combination of Focal (Lin et al., 2017) and Dice (Sudre et al., 2017) losses to supervise the predicted 3D masks effectively.
+Optimal matching, $\mathbf{d}_{\text{match}}$: this function measures the pairwise distance between the 3D grounding results and the ground-truth values, and is used for matching.
+It is implemented similarly to $\mathcal{L}_{\text{ground}}$ but with different loss weights.
+The model always predicts the maximum number of instances (256) but avoids false positive detections by matching the unused instance to a special no-match text token (Jain et al., 2021).
+The maximum of 256 was not a major limitation in practice, but could be increased.
+3.2.2 Per-Pixel Losses
+While grounding loss alone is sufficient for stable pretraining (Table 4), LIFT-GS benefits from joint training with additional photometric and feature losses for faster convergence.
+These losses are only used during pretraining, not for finetuning with 3D annotations.
+Reconstruction loss, $\mathcal{L}_{\text{RGB}}$: supervises photometric reconstruction using $L_{1}$ and SSIM losses (Hong et al., 2024; Zhu et al., 2023b; Wang et al., 2004)
 $\mathcal{L}_{\text{RGB}}=\lambda_{1}\mathcal{L}_{1}(I,\tilde{I})+\lambda_{2}\mathcal{L}_{\textbf{SSIM}}(I,\tilde{I})$ (6)
-Grounding losses: The grounding losses, $\mathcal{L}{\text{mask}}$, measure the prediction quality between the rendered 2D masks, $\tilde{\mathbf{M}}_{\text{2D}}$, and the ground-truth masks, $\mathbf{M}_{\text{2D}}$.
-Since the predicted and ground-truth masks differ in number, we employ Hungarian matching to compute the set-to-set loss.
-$\sigma(i)$ is the index of predicted masks matching to $i$-th ground truth mask according to the matching distance $\mathbf{d}_{\text{match}}$.
-$\sigma(i)=\arg\min_{j}\mathbf{d}_{\text{match}}(\tilde{\mathbf{M}},\mathbf{M}_{i},\mathbf{C})$ (3)
-The mask shape is supervised with $\mathcal{L}_{\text{mask}}$, the combination of Focal (Lin et al., 2017) and Dice (Sudre et al., 2017) loss used in SAM (Kirillov et al., 2023):
-$\mathcal{L}_{\text{ground}}\!=\!\frac{1}{K}\sum_{i}^{K}\lambda_{\text{3}}\mathcal{L}_{\text{mask}}(\tilde{\mathbf{M}}_{\text{2D}}^{\sigma(i)},\mathbf{M}_{\text{2D}}^{i})\!+\!\lambda_{\text{4}}\mathcal{L}_{\text{CE}}(\mathbf{C}_{\sigma(i)},i)$ (4)
-$\mathcal{L}_{\text{CE}}$ is the correspondence loss, encouraging correct correspondences between predicted masks and language tokens.
-Specifically, the ground-truth mask set $\mathbf{M}_{\text{2D}}$ consists of $K$ binary masks, each corresponding to a span $\mathbf{Q}_{k}\subseteq\mathbf{Q}$ of language query embeddings $\mathbf{Q}$.
-$\mathbf{C}_{\sigma(i)}$ represents the correspondence probability of the $\sigma(i)$-th mask proposal across spans of the $|Q|$ language embeddings.
-We optimize this using a cross-entropy loss.
-$\mathcal{L}_{\text{CE}}(\mathcal{C}_{\sigma(i)},i)=-\log\frac{\exp(\mathbf{C}_{\sigma(i),i})}{\sum_{j}^{K}\exp(\mathbf{C}_{\sigma(i),j})}$ (5)
-Feature loss: $\mathcal{L}_{\text{feat}}$ is the feature rendering loss which measures the difference between rendered feature map $\tilde{\mathbf{F}}_{\text{2D}}$ and ground-truth feature map $\mathbf{F}_{\text{2D}}$.
-As in (Zhu et al., 2023b), LIFT-GS uses a contrastive loss.
+Feature loss, $\mathcal{L}_{\text{feat}}$: uses CLIP-style contrastive regularization to align rendered features $\tilde{\mathbf{F}}_{\text{2D}}$ with ground-truth features $\mathbf{F}_{\text{2D}}$ as in (Zhu et al., 2023b):
 $\mathcal{L}_{\text{feat}}=\frac{1}{H\times W}\sum_{u,v}^{H,W}-\log\frac{\exp(\tilde{\mathbf{f}}_{(u,v)}\cdot\mathbf{f}_{k})}{\sum_{j}\exp(\tilde{\mathbf{f}}_{(u,v)}\cdot\mathbf{f}_{j})}$ (7)
-$\tilde{\mathbf{f}}{(u,v)}$ is the feature vector of $\tilde{\mathbf{F}}_{\text{2D}}$ at location $(u,v)$.
-${\mathbf{f}}_{j}$ denotes a batch of unique feature vectors from $\mathbf{F}$, and $\mathbf{f}k$ is the ground-truth feature vector corresponding to $\tilde{\mathbf{f}}{(u,v)}$.
+where $\tilde{\mathbf{f}}_{(u,v)}$ is the rendered feature at pixel $(u,v)$ from $\tilde{\mathbf{F}}_{\text{2D}}$, $\mathbf{f}_{k}$ is the corresponding ground-truth feature, and $\mathbf{f}_{j}$ represents the batch of unique features.
+3.3 Architecture
+LIFT-GS is network-agnostic with minimal architectural constraints.
+The design can be applied to other architectures, with constraints only arising from the specific losses used.
+This section describes the network used in our experiments, shown in Figure 5.
+3.3.1 Grounding Decoder: (see Grounding Losses)
+The grounding decoder is a transformer based on MaskFormer (Cheng et al., 2021b) that predicts the correspondence matrix $\mathbf{C}\in\mathbb{R}^{m\times|Q|}$ and 3D Gaussian masks $\mathbf{M}\in\mathbb{R}^{N\times m}$.
+LIFT-GS uses the mask decoder from UniVLG (Jain et al., 2025) with minimal modifications.
+The transformer takes Gaussian features $\mathbf{G}$ and language query embeddings $\mathbf{Q}$ as inputs, using a set of learnable tokens as 3D mask proposals.
+Cross-attention is computed between the learnable tokens and both the language and Gaussian tokens.
+After extensive information exchange within the transformer, $\mathbf{M}$ (or $\mathbf{C}$) is computed as dot products between mask proposal tokens and Gaussian tokens (or mask proposal tokens and language tokens, respectively).
+The detailed LIFT-GS architecture and hyperparameter settings used in our experiments are provided in Appendix A.
+3.3.2 Gaussian Decoder Head: (see Per-Pixel Losses)
+LIFT-GS predicts $\mathbf{G}\in\mathbb{R}^{N\times F}$ using a learned pointwise MLP applied to the point cloud encoder outputs.
+While the number of predicted Gaussians $|G|$ can differ from input points $|P|$, we set $|G|=|P|$ with a bijective mapping for consistency with point cloud evaluation tasks.
+Notably, the Gaussian decoder does not require direct supervision–3D Gaussians can be treated as latent variables for 3D VLG, as shown in Table 4 when per-pixel losses are disabled.
+3.3.3 Input Encoders:
+Pointcloud Encoder: Tokenizes RGB point clouds of shape $|P|{\times}6$ (xyz + RGB) using a sparse convolutional UNet (Contributors, 2022), following PonderV2 (Zhu et al., 2023b).
+Weights are randomly initialized and learned.
+Text Encoder: LIFT-GS uses CLIP text embeddings, which remain frozen during training.
+These encoders represent common choices in 3D VLG. Stronger architectural choices like transformer-based pointcloud encoders or different text embeddings would likely improve performance.
 3.4 SAM-CLIP 2D Pseudo-Label
-Although LIFT-GS doesn’t need 3D annotations, getting high-quality 2D supervision remains challenging.
-We show how leveraging 2D foundation models for pseudo-label generation enables reasonable zero-shot performance and significantly enhances downstream tasks after fine-tuning.
-As shown in Figure 3, we leverage 2D foundation models to generate pseudo-labels, including pseudo language queries and corresponding 2D masks.
-For each image, we apply SAM (Kirillov et al., 2023) to obtain segmentation masks.
-For each segmented region, we extract a CLIP (Radford et al., 2021) image embedding as the pseudo language query embedding.
-Since CLIP’s text and image embeddings share the same feature space, LIFT-GS can take text embeddings as input during inference.
-During pretraining, we concatenate these CLIP embeddings to form $\mathbf{Q}$ and construct $\mathbf{C}$, using the corresponding 2D masks as ground truth.
+While LIFT-GS eliminates the need for 3D annotations, obtaining high-quality 2D supervision remains challenging.
+We show one way in which 2D foundation models can generate pseudo-labels that enable reasonable zero-shot performance and significantly enhance downstream fine-tuning.
+As shown in Figure 4, we generate pseudo-labels using SAM (Kirillov et al., 2023) and CLIP (Radford et al., 2021).
+For each image, SAM provides segmentation masks, and for each segmented region, we extract CLIP image embeddings as pseudo language query embeddings.
+Since CLIP’s text and image embeddings share the same feature space, LIFT-GS can use text embeddings during inference.
+We concatenate these CLIP embeddings to form $\mathbf{Q}$ and construct $\mathbf{C}$ s.t.
+each instance maps to exactly one query token.
 Figure 6: Zero-Shot 3D Segmentation.
@@
 Language queries include both high-level abstract concepts (e.g., white) and detailed descriptions (e.g., black cabinet near the wall).
-Trained with 2D pseudo-labels, LIFT-GS can perform zero-shot 3D grounding using real text queries without fine-tuning, as shown in Figure 5.
-However, the zero-shot model suffers from low accuracy and struggles with complex expressions, a common limitation of CLIP-based methods that function as bag-of-words models (Yuksekgonul et al., 2023).
-Future improvements in pseudo-labeling, such as captioning (Meta AI, 2024) and 2D language grounding models (Liu et al., 2023), could alleviate the need for fine-tuning altogether.
-In this work, we focus on fine-tuning.
-Overall, LIFT-GS shows that even simple pseudo-labeling strategies can be effectively distilled into 3D models.
-Our experiments show that pretraining with 2D pseudo-labels substantially boost performance on downstream tasks.
-3.4 Architecture
-LIFT-GS  is network-agnostic, imposing few architectural constraints and being readily adapted to other architectures. We describe the network used in experiments, as in Figure 4.
-Encoder: Predicting 3DGS from Point Cloud
-Encoder takes point clouds as input and predicts 3D Gaussian splatting (without masks) and per-point features as latent variables.
-Generally, the number of predicted Gaussians $|G|$ can differ from the number of input points $|P|$.
-However, our focus is on demonstrating the effectiveness of differentiable rendering for structured prediction tasks like Promptable 3D segmentation.
-So in our experiments, we set $|G|=|P|$ with a bijective mapping to ensure consistency with point cloud-based evaluation tasks.
-For comparison to prior work (Zhu et al., 2023b) , LIFT-GS uses a SparseConv UNet (Contributors, 2022) as the Encoder Backbone to get per-point features, and uses an MLP as Gaussian Head to regress Gaussian parameters of $\mathbf{G}$ with features $\mathbf{F}$.
-Decoder: Mask Decoder for 3D VLG
-The Decoder takes per-point features from the Encoder and language embeddings $\mathbf{Q}$ as input to predict the 3D mask $\mathbf{M}\in\mathbb{R}^{N\times m}$ and the correspondence matrix $\mathbf{C}\in\mathbb{R}^{m\times|Q|}$.
-The predicted $\mathbf{M}$ serves as a renderable feature of the 3DGS $\mathbf{G}$ and is subsequently used to generate 2D masks.
-While this work focuses on 3D vision-language grounding (3D VLG), the proposed framework is a general pipeline that can be extended to other 3D tasks with renderable outputs.
-We adopt a 3D variant of the Transformer-based mask decoder in MaskFormer (Cheng et al., 2021a; Jain et al., 2025), inspired by the success of MDETR (Kamath et al., 2021) and BUTD (Jain et al., 2021).
-As illustrated in Figure 4, the 3D Mask Decoder Transformer takes three inputs: visual tokens from per-point features, language tokens from $\mathbf{Q}$, and $m$ learnable mask proposal tokens.
-It outputs $m$ predicted mask embeddings and a correspondence matrix $\mathbf{C}$.
-The 3D masks $\mathbf{M}$ are computed as the dot product between visual tokens and output mask embeddings.
-These similarity scores are then differentiably rendered into mask images $\tilde{\mathbf{M}}_{\text{2D}}$.
-4 Experiments
+With 2D pseudo-labels, LIFT-GS performs zero-shot 3D grounding using real text queries without fine-tuning (Figure 6).
+However, zero-shot performance suffers from low accuracy and struggles with complex expressions – a common limitation of CLIP-based methods that function as bag-of-words models (Yuksekgonul et al., 2023).
+Future improvements in pseudo-labeling, such as better captioning (Meta AI, 2024) and 2D language grounding models (Liu et al., 2023), could reduce reliance on fine-tuning.
+LIFT-GS demonstrates that even simple pseudo-labeling strategies can be effectively distilled into 3D models.
+In the experiments below, pretraining with 2D pseudo-labels substantially improves downstream task performance.
+Fine-tuning data scaling results are consistent with established transfer learning “scaling laws” (Hernandez et al., 2021).
 Table 1: Open-Vocabulary 3D Instance Segmentation. We evaluate our model on ScanNet200 by using category names as text queries and compare it against SOTA models.
@@
 $\Delta$ $+3.2\uparrow$ $+5.1\uparrow$ $+4.3\uparrow$
-In this section, we first provide training details and demonstrate how pretraining significantly enhances downstream task performance through a set of carefully designed ablations.
-Additionally, we uncover several intriguing findings by scaling pretraining and fine-tuning data, as well as exploring the impact of different 2D foundation models.
+4 Experiments
+Although supervised via 2D loss, LIFT-GS is a fully 3D model that explicitly outputs 3D masks.
+Beyond the zero-shot setting using 2D pseudo-labels (Figure 6), LIFT-GS can be readily fine-tuned with 3D annotation data using 3D losses for significantly improved performance.
+To enhance practical applicability, we focus on fine-tuning with 3D annotations and demonstrate how 2D model distillation boosts performance.
+In this section, we first provide training details and show how pretraining significantly improves downstream task performance through a series of carefully designed ablations.
+We also reveal several insights from scaling the amount of pretraining and fine-tuning data, and explore the impact of different 2D foundation models.
 4.1 Training Details
-We provide details below with more details in Appendix.
-Training Datasets We use ScanNet (Dai et al., 2017) as the primary dataset for downstream task fine-tuning and evaluation, as its annotations form the basis for established benchmarks.
-Our method’s flexibility enables training on diverse 3D datasets without requiring language annotations or instance masks.
-For pretraining, we primarily use ScanNet (Dai et al., 2017) and ScanNet++(Yeshwanth et al., 2023) and also explore Taskonomy(Zamir et al., 2018) and Aria Synthetic (Somasundaram et al., 2023).
-Architecture Our pipeline imposes no constraints on the 3D architecture or inductive biases, allowing flexibility in model selection.
-For experiments, we use SparseConv UNet as the encoder backbone, a widely adopted architecture for point cloud processing.
-The Mask Decoder is an 8-layer Transformer decoder with a hidden state size of 512.
-Training Details We train the Encoder and Decoder jointly using a batch size of 32 for 76k steps on 32 A100 GPUs.
-We use AdamW (Loshchilov & Hutter, 2017) optimizer with a learning rate of 1e-4 and weight decay of 1e-4.
+We provide details below with more details in the Appendix.
+Datasets We use ScanNet (Dai et al., 2017) as the primary dataset for downstream task fine-tuning and evaluation, as its annotations form the basis for established benchmarks.
+We primarily pretrain using ScanNet (Dai et al., 2017) for comparison to other methods.
+LIFT-GS enables training on diverse unlabeled 3D datasets, and we show additional pretraining scaling experiments using ScanNet++(Yeshwanth et al., 2023), Taskonomy (Zamir et al., 2018) and Aria Synthetic (Somasundaram et al., 2023).
+Architecture LIFT-GS method imposes minimal architectural constraints.
+In the following experiments, backbones consist of a text encoder (frozen CLIP-L), point cloud encoder (Sparse 3D UNet (Çiçek et al., 2016)), and grounding decoder (8-layer Transformer, hidden size 512 following (Jain et al., 2025)), and an MLP for the Gaussian decoder.
+Training Models are trained end-to-end for 76k steps with batch size 32 on 32 A100s using AdamW (Loshchilov & Hutter, 2017) (lr=1e-4, weight decay=1e-4).
+Point clouds are voxel-downsampled to 5cm for an average of  50k points, and we render masks at resolution $512{\times}512$ to ensure small masks are captured.
+Complete implementation details are provided in the appendix.
 4.2 Evaluation on 3D Vision-Language Grounding
@@
 Since LIFT-GS outputs masks instead of axis-aligned bounding boxes, we derive bounding boxes by extracting the extreme corner points from the point cloud within the predicted masks.
-LIFT-GS is designed to be practical.
-To ensure a more realistic evaluation for embodied applications, we introduce two modifications: (1) we predict 3D masks without assuming the availability of ground-truth 3D bounding boxes, and (2) we utilize sensor point clouds (Sensor PC) from RGB-D scans instead of mesh-derived point clouds (Mesh PC).
-Existing methods are typically trained on mesh-derived point clouds (Mesh PC), requiring creating scene meshes from RGB-D scans, processing, and sampling point clouds.
-They generally rely on segments derived from clustering centers generated during human annotation.
-Since such information is unavailable in real-world applications, we instead use raw sensor point clouds (Sensor PC) from RGB-D scans.
-This setting is inherently more challenging, as reflected in the significant performance drop of BUTD-DETR (Jain et al., 2021) when transitioning from Mesh PC to Sensor PC (Table 2), consistent with findings in (Jain et al., 2024).
+LIFT-GS is designed to be practical and we evaluate it using the “real-world” settings used in more recent 3D VLG work where (1) we predict 3D masks without assuming known ground-truth 3D bounding boxes, and (2) we utilize sensor point clouds (Sensor PC) from RGB-D scans instead of using mesh-derived point clouds that leak label information (Mesh PC).
+This realistic setting is more challenging, as reflected in the significant performance drop of BUTD-DETR (Jain et al., 2021) when transitioning from Mesh PC to Sensor PC (Table 2), consistent with findings in (Jain et al., 2024).
+A more complete comparison of these settings is provided in (Jain et al., 2021, 2025; Arnaud et al., 2025).
 Baselines We compare LIFT-GS against the state-of-the-art two-stage methods, 3D-VisTA (Zhu et al., 2023c) and PQ3D (Zhu et al., 2024), as well as the SOTA single-stage method, BUTD-DETR (Jain et al., 2021).
@@
 Notably, LIFT-GS outperforms 3D-VisTA in Acc@$25$, despite 3D-VisTA being a two-stage method with bounding box proposals from Mask3D using Mesh PC.
-4.3 Ablation and Analyses
+4.3 Pretraining Ablations
 We conduct an in-depth analysis of the proposed method through a series of ablation and scaling experiments.
@@
 Additionally, we report the higher accuracy threshold Acc@$0.75$.
-Compare to SOTA pretraining methods We compare to Ponder-v2 (Zhu et al., 2023b), a state-of-the-art method for point cloud pretraining.
-Ponder-v2 takes point clouds as input and predicts voxel grids, which are then used to render 2D images via NeRF (specifically NeuS, which additionally models surfaces (Wang et al., 2021)).
-Pretraining is supervised by 2D photometric loss and CLIP features, where features are computed based on per-pixel text labels for the mask (GT category labels and masks for each instance).
-We found that the official Ponder-v2 checkpoint fails to improve performance on 3D referential grounding, likely due to the limited size and diversity of its per-pixel text labels.
-Therefore, we retrain the Ponder-v2 while using the SAM-CLIP pseudolabels described in Section 3.3.
-Training on the pseudolabels (indicated by Ponder-v2$\dagger$ in Table 3) significantly improves Acc@$0.25$ +4.5% vs.
-the official Ponder-v2, and +3.2% vs.
-training from scratch.
-Since both approaches use render-supervision, this experiment underscores the advantages of pseudolabel distillation in data-scarce regimes.
-Moreover our method does not rely on human-annotated per-pixel text labels, making it highly scalable to other datasets, as shown in Table 6.
+Compare to SOTA pretraining methods We compare against PonderV2 (Zhu et al., 2023b), a state-of-the-art point cloud pretraining method that also uses render-supervision.
+Since the official PonderV2 relies on limited human-annotated text labels, we retrain it using our SAM-CLIP pseudo-labels for fair comparison (PonderV2$\dagger$ in Table 3).
+This demonstrates the value of distillation, improving the performance over using GT labels (45.4% vs 40.9% Acc@0.25).
+Moreover, LIFT-GS substantially outperforms PonderV2$\dagger$ (47.5% vs 45.4% Acc@0.25), underscoring the impact of multimodal decoder architectures enabled by the LIFT-GS render-supervised formulation.
 Table 4: Loss Ablation.
@@
 - 47.53 33.75 13.49
-Loss Ablation Existing pretraining pipelines primarily focus on the encoder (Zhu et al., 2023b; Banani et al., 2021), whereas the render-supervised formulation can pretrain the entire architecture in a unified manner.
-To assess the impact of the grounding loss, which applies only to the decoder, we compare models pretrained with different losses in Table 4.
-A model trained with $\mathcal{L}_{\text{ground}}$ alone (row 2) still substantially improves downstream task performance, performing only slightly worse than the model trained with the full loss (row 5), demonstrating the effectiveness of the grounding loss.
+Loss Ablation Existing pretraining pipelines primarily focus on the encoder (Zhu et al., 2023b; Banani et al., 2021), whereas the render-supervised formulation can pretrain the entire architecture in a unified manner using the grounding loss.
+We find that grounding loss alone can be used to pretrain the model end-to-end in Table 4.
+A model trained with $\mathcal{L}_{\text{ground}}$ alone (row 2) substantially improves downstream task performance, performing only slightly worse than the model trained with all losses (row 5).
 Furthermore, comparing models with and without $\mathcal{L}{\text{ground}}$ (row 5 vs.
@@
 4.4 Data Scaling
-Finetuning Data Scaling We limit the fine-tuning data size to $0.1$, $0.2$, and $0.5$ of the full dataset for 3D referential grounding to analyze performance variation with different fine-tuning ratios.
-The results are presented in Figure 6 and Table 5, with additional findings provided in Appendix A.3.
-Intriguingly, we find that LIFT-GS “pretraining effectively multiplies the fine-tuning dataset” (Hernandez et al., 2021) by a constant factor (roughly 2x), across varying amounts of fine-tuning data.
-For instance, the pretrained model fine-tuned on $20\%$ of the data achieves performance comparable to training from scratch with $50\%$ of the data.
-The benefits of pretraining are even more pronounced at higher IoU thresholds.
-Overall, a pretrained model can achieve the same performance as a model trained from scratch on all the data using only $30{-}40\%$ of the fine-tuning data.
-This observation matches empirical data scaling laws for pretraining in other modalities (Hernandez et al., 2021).
-And interestingly, the fact that the coefficient does not show diminishing returns, even when we use all available fine-tuning data, which underscores that 3D VLG models are currently operating in the data scarce regime.
+LIFT-GS exhibits strong scaling properties that reveal 3D VLG operates in a severely data-scarce regime.
+Finetuning Data Scaling We observe that pretraining effectively ”multiplies” the fine-tuning dataset by approximately 2x.
+As shown in Figure 7 and Table 5, a pretrained model using $50\%$ of fine-tuning data matches the performance of training from scratch with $100\%$ data.
+This scaling coefficient remains constant across different data amounts ($10\%$, $20\%$, and $50\%$) without diminishing–matching empirical scaling laws from other modalities (Hernandez et al., 2021).
+The benefits are most pronounced at higher IoU thresholds, where a pretrained model achieves scratch-level performance using only $30{-}40\%$ of the fine-tuning data.
+Additional results on Instance Segmentation are provided in Appendix A.3.
 Figure 7: Fine-tune Data Scaling. We show how Grounding Accuracy changes with increasing Data Ratio from 0.1 to 1.0.
-Pretraining Data Scaling A key advantage of our framework is its ability to pretrain on large-scale RGB-D scans without requiring 3D or language annotations, ensuring scalability.
-We scale the pretraining dataset and analyze its impact on fine-tuning performance, as shown in Table 6.
-[…]
-Since our method doesn’t need 3D masks or language annotations, expanding the pretraining dataset with real-world scans is fairly easy.
-We anticipate continuous improvements as larger-scale data collections become available.
+Pretraining Data Scaling Expanding pretraining data consistently improves downstream performance (Table 6).
+Adding ScanNet++ yields notable gains (+0.8%), while incorporating Taskonomy and Aria Synthetic provides additional improvements despite distribution differences (likely due to the mesh reconstruction quality in Taskonomy).
+Comparing the scaling results from pretraining and fine-tuning demonstrates the strong data efficiency of LIFT-GS.
+As shown in Table 6, adding ScanNet++ ( 30% of the data used for pretraining) yields a performance gain equivalent to adding 15% more ScanNet fine-tuning data with 3D annotations, based on the curve in Figure 7.
+This indicates an effective transfer ratio of roughly 1:2—i.e., collecting twice as many raw videos provided improvements comparable to building a fully annotated 3D SFT dataset.
+Therefore, pretraining on ScanNet++ is not only highly effective but also cost-efficient, especially considering that annotating 3D referential grounding data requires significantly more effort than collecting raw videos alone.
+Data Scarcity Implications The consistent 2× multiplier without saturation, combined with continued gains from more pretraining data, strongly suggests that current 3D VLG models are severely limited by data availability–opening a path for future 3D VLG improvements through scaling alternate sources (such as 2D foundation models).
 Table 6: Pretraining on OOD data. Adding more pretraining data from ScanNet++ improves performance. Taskonomy and Arial helped less than ScanNet++, likely due to distribution difference.
@@
 Replacing them with smaller models, MobileSAM (Zhang et al., 2023a)(ViT-tiny)  and CLIP-B, leads to a noticeable performance drop, especially at higher accuracy thresholds.
-This suggests that our model benefits from advancements in 2D foundation models, highlighting the importance of stronger 2D models.
+This suggests that render-supervised distillation directly benefit from advancements in 2D foundation models.
 Table 7: 2D Foundation Model Exploration.
@@
 SAM-H + LLAMA-Caption 47.50 32.78 13.25
-Captions from VLMs We explore an alternative pseudo-labeling strategy by combining SAM with the LLAMA-V model (details in Appendix).
-After segmenting objects in 2D images using SAM, we prompt LLAMA-V to describe the segmented regions and encode these descriptions into CLIP embeddings.
-Pretraining with this pseudo-labeling method achieves performance comparable to our original pipeline with SAM-H and CLIP-L, highlighting the flexibility and extensibility of our framework.
-We believe text-based captions hold significant potential for future research, and our pipeline is well-positioned to benefit from advancements in this area.
+Captions from LMMs Table 7 shows results using large LMMs instead of CLIP to generate queries (LLAMA-3V + SAM grounding, details in Appendix).
+After segmenting objects in 2D images using SAM, we prompt LLAMA-V to describe the segmented regions.
+Pretraining with these captions achieves performance comparable to our original pipeline with SAM-H and CLIP-L.
+As LMMs continue to improve, we believe text-based captions hold significant potential for future research, and the approach highlighted in the paper is positioned to benefit from LMM improvements.
 5 Conclusion
-We present LIFT-GS, a model-agnostic pipeline for training 3D VLG models without 3D supervision by leveraging Gaussian Splatting and 2D foundation models.
-Beyond enabling zero-shot 3D grounding, LIFT-GS significantly improves downstream performance when fine-tuned with limited 3D data.
-It also scales effectively, benefiting from larger datasets and stronger 2D foundation models, suggesting continuous improvements with increased data.
-Limitations and future directions are discussed in the Appendix.
+LIFT-GS tackles data scarcity that limits 3D VLG by introducing render-supervised distillation from 2D VLM models.
+By training 3D models using only 2D supervision from models like SAM and CLIP, LIFT-GS achieves state-of-the-art performance for 3D VLG.
+Our findings, including consistent 2x data multiplication effects, reveal that 3D grounding currently operates with substantial data limitations.
+LIFT-GS circumvents a key 3D annotation bottleneck by introducing a scalable training approach that benefits directly from advancements in frontier multimodal language models.
+It offers a practical technique to leverage progress in 2D to accelerate the development of other data-scarce capabilities essential for robotics, AR/VR, and embodied AI.
+Acknowledgment
+This work is done during Ang Cao’s internship at Meta. We thank Andrew Owens, Andrea Vedaldi, Stella Yu, Ziyang Chen, Yiming Dou, Xuanchen Lu for their helpful discussion and feedback.
+Impact Statement
+We propose a method for training 3D models without 3D supervision, advancing 3D vision-language research.
+Our approach significantly improves 3D referential grounding, a key task for robotics, embodied AI, and AR/VR applications.
+The resulting model enables agents to precisely localize objects from language inputs, bridging high-level reasoning with real-world actions.
+As our model is distilled from 2D foundation models, it may inherit their biases. However, since our primary task is grounding, it is unlikely to introduce significant aesthetic biases.
 Appendix A More Details